### v1

# Personalized Item Embeddings in Federated Multimodal Recommendation

Abstract.
Federated recommendation systems play a crucial role in protecting user privacy. However, existing methods primarily rely on ID-based item embeddings, overlooking the rich multimodal information of items. To address this limitation, we propose a novel Federated Multimodal Recommendation System called FedMR. FedMR leverages a foundation model on the server side to encode multimodal data, such as images and text, associated with items. To tackle the challenge of data heterogeneity caused by varying user preferences, FedMR introduces a Mixing Feature Fusion Module on the client. This module dynamically adjusts the weights of different fusion strategies based on user interaction history, generating personalized item embeddings that capture fine-grained user preferences. FedMR is compatible with existing ID-based federated recommendation systems, improving their performances without modifying the original framework. Our experiments on four real-world multimodal recommendation datasets demonstrate the effectiveness of FedMR. Our code is available at https://anonymous.4open.science/r/FedMR.
Federated Recommendation, Multimodal Learning, Personalization
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Collaborative filteringccs: Information systems Personalizationccs: Information systems Combination, fusion and federated search
1. Introduction
Under privacy protection regulations, such as GDPR (Voigt and Von dem Bussche, 2017), that emphasize user data privacy, Federated Recommendation Systems (FedRec) have become an essential technology in modern life (Sun et al., 2024). However, in these systems, since each client typically only has access to a single user’s interaction history, data sparsity is a common issue that degrades recommendation performance (Li and Long, 2024). Modeling item features as shared information across clients can effectively mitigate the impact of data sparsity. Many studies (Zhang et al., 2023c; Li et al., 2023a, 2024b) have shown that personalized item feature representations can significantly improve recommendation results. Despite items often containing rich multimodal information (Zhou et al., 2023b), such as such as text and images, most existing FedRec methods (Zhang et al., 2023d, 2024d, 2024f) still rely mainly on ID embeddings, underutilizing the potential of multimodal data. Thus, generating personalized item embeddings using multimodal data while preserving user privacy remains a significant challenge.
As a potential solution, Multimodal Recommendation Systems (MMRec) integrate data from different modalities, significantly enhancing recommendation performance (He and McAuley, 2016; Wei et al., 2019; Zhou et al., 2023a). Unlike single-modal systems, MMRec can capture hidden relationships and complementary information between various modalities, leading to a better understanding of user preferences and item characteristics (Zhou et al., 2023b). For example, MMRec methods (Wei et al., 2019; Wang et al., 2021; Tao et al., 2022) combine video frames, audio tracks, and item descriptions when users interact with online platforms, providing a more accurate representation of user preferences. To effectively handle multimodal data, MMRec often uses Foundation Models (FM) pretrained on large-scale datasets, which possess powerful feature extraction and representation capabilities (Bommasani et al., 2021). By applying FMs, MMRec can efficiently extract and integrate information from multiple modalities, capturing commonalities and complementary relationships between them (Radford et al., 2021a; Cui et al., 2022; Geng et al., 2022, 2023; Wu et al., 2024). However, these centralized MMRec methods rely on servers to aggregate user data, digital activities, and preferences to train accurate models (Liu et al., 2024), which raises privacy concerns (Chai et al., 2020; Li et al., 2023a).
To address these challenges, Federated Multimodal Recommendation Systems (FedMMRec) have emerged, which integrates MMRec and introduces the capability of processing multimodal information into FedRec. By utilizing multimodal data and user-specific information, such as profiles and interaction history, locally on each client to generate personalized recommendations and aggregating model parameters at the server, FedMMRec enhances recommendation performance while preserving user privacy (Feng et al., 2024). FedMMRec is still in the early stages of research, and related studies are limited. FedMMR (Wu et al., 2024) adopts a model-splitting strategy where the computationally intensive components are handled by the server, while clients perform lightweight collaborative filtering. AMMFRS (Feng et al., 2024) introduces an attention mechanism on the client side to perform weighted fusion of multimodal data, effectively distinguishing and utilizing the contributions of different modalities. These methods not only protect user privacy but also effectively integrate multimodal features to improve recommendation accuracy and personalization. However, they rely on fixed multimodal fusion strategies, neglecting the personalized representation of item features on clients, which limits their ability to adapt to diverse user needs. Although AMMFRS uses an attention mechanism to distinguish the importance of different modalities, it lacks adaptability to the diversity of user data, resulting in limited flexibility and accuracy of fusion strategies (Niu et al., 2021).
Main Contribution. To integrate multimodal item information into FedRec while ensuring user privacy, we propose a Federated Multimodal Recommendation system (FedMR) based on foundation models. FedMR uses a foundation model on the server side to encode multimodal item data (e.g., images, text) into feature embeddings, and constructs a global item ID embedding to capture structural features among items. On each client, FedMR introduces a Mixing Feature Fusion Module that dynamically adjusts fusion strategies based on user interaction history, combining multimodal and ID embeddings to generate personalized item embeddings. This approach provides more personalized recommendations. Thus, FedMR seamlessly integrates multimodal information with existing ID-based FedRec methods, improving the flexibility and accuracy of recommendations without altering the existing framework, and effectively handles data heterogeneity across clients.
The main contributions of this work are summarized as follows:
  • We propose FedMR, a federated multimodal recommendation system using foundation models, which efficiently leverages multimodal data while preserving user privacy;
  • A Mixing Feature Fusion Module is designed to dynamically adjusts fusion strategies based on user interaction histories to generate personalized item embeddings, enhancing personalization and flexibility in recommendations;
  • FedMR is compatible with existing ID-based federated recommendation methods, improving recommendation performance without modifying the existing framework;
  • Extensive experiments are conducted on four real-world multimodal recommendation datasets to validate the effectiveness of our proposed method FedMR;
  • We provide the source code to support the reproducibility.
2. Related Work
2.1. Federated Recommendation Systems
Federated recommendation systems (FedRec) aim to provide personalized recommendations while ensuring user privacy and data security (Yang et al., 2020). In recent years, FedRec has made significant progress in areas such as rating prediction (Lin et al., 2020a, b; Liang et al., 2021; Lin et al., 2022; Zhang et al., 2023d; Li et al., 2024a), top-k recommendation (Ammad-Ud-Din et al., 2019; Perifanis and Efraimidis, 2022; Luo et al., 2022; Zhang et al., 2023b; Dong et al., 2022; Li et al., 2022a; Qu et al., 2023), and addressing the cold-start problem (Liu et al., 2021b; Meihan et al., 2022; Wahab et al., 2022; Zhang et al., 2024c, d). However, most existing FedRec methods (Huang et al., 2020; Flanagan et al., 2021; Wu et al., 2022; Zhu et al., 2022; Imran et al., 2023; Yuan et al., 2023a, 2024) rely heavily on ID embeddings to model items, overlooking the rich multimodal information (e.g., titles and covers) that items may have. AMMFRS (Feng et al., 2024) and FedMMR (Wu et al., 2024) apply multimodal learning (Baltrušaitis et al., 2018) to FedRec. AMMFRS uses an attention mechanism (Niu et al., 2021) to weight and fuse multimodal data, emphasizing the contribution of different features. In contrast, FedMMR assigns the computationally heavy multimodal learning tasks to the server and introduces a secure cross-client convolution strategy to capture high-order interaction information using secret sharing techniques while preserving privacy. Unlike methods that focus only on personalizing user information, PFedRec (Zhang et al., 2023c) proposes a dual personalization mechanism that optimizes personalization for both users and items, integrating it into existing FedRec frameworks. Building on this, FedRAP (Li et al., 2023a) introduces additive personalization, which refines the modeling of user and item features. FedDAE (Li et al., 2024b) employs a dual-encoder variational autoencoder to adaptively fuse shared and personalized item knowledge through a personalized gating network. While these methods achieve personalized item embeddings, they still mainly rely on ID embeddings, failing to fully utilize the rich multimodal information of items.
2.2. Multimodal Recommendation Systems
Unlike traditional Recommendation Systems (RecSys) that rely primarily on user-item interaction data (such as clicks or ratings) to capture user behavior (Yu et al., 2023), Multimodal Recommendation systems (MMRec) aim to improve the accuracy of recommendations and enhance user experience by utilizing multiple modalities of data , such as text, images, audio, and video (Zhou et al., 2023b). As unique identifiers for items, item IDs not only contain semantic information about the items but also capture the characteristics of neighboring items (Yuan et al., 2023b; Liu et al., 2023). Therefore, in MMRec, item IDs are often used as auxiliary information and combined with multimodal data to enhance item representations(Wang et al., 2024; Zhang et al., 2024e). This integration helps capture a more comprehensive understanding of users’ interests and preferences, ultimately improving the quality of recommendations. A typical MMRec training process generally involves feature extraction and modality fusion (He and McAuley, 2016; Wei et al., 2019; Wang et al., 2021; Tao et al., 2022). Feature extraction (Guo et al., 2019) aims to obtain representative features from multiple modalities to better capture the attributes and preferences of users and items. Modality fusion (Gao et al., 2020) focuses on effectively integrating data from different modalities to enhance the overall performance of the MMRec and the relevance of its results. The recommendation models (Tao et al., 2020; Wu et al., 2021; Zhou and Shen, 2023) combine multimodal data with user-item interaction data to predict user preferences and provide personalized recommendations.
2.3. Foundation Models
Foundation Models (FM) are often pre-trained on large-scale unlabeled datasets for general purposes, that can be adapted to a variety of downstream tasks through appropriate fine-tuning, such as task-specific training (Bommasani et al., 2021). This makes them highly generalizable and adaptable across different applications. They are then adapted to specific recommendation tasks through methods such as representation transfer (Du et al., 2020; Liu et al., 2021a, 2022; Wei et al., 2023), model fine-tuning (Zhang et al., 2021; Li et al., 2022b, 2023b; Sun et al., 2023; Wang et al., 2023), adapter tuning (Hou et al., 2022; Bian et al., 2023; Geng et al., 2023; Fu et al., 2024), and prompt tuning (Shin et al., 2020; Li and Liang, 2021; Zhou et al., 2022; Zhang and Wang, 2023; Xu et al., 2024). Some studies have explored the use of FM to enhance the content-awareness of MMRec, improving content understanding and generation by leveraging alignment and generation techniques across multiple modalities. CLIP (Radford et al., 2021a; Wu et al., 2024) uses contrastive learning (Le-Khac et al., 2020) between language and images to obtain multimodal representations, capturing the matching relationships between users and items. M6-Rec (Cui et al., 2022) builds a recommendation system based on large-scale multimodal pre-trained models specifically designed for the e-commerce domain. P5 (Geng et al., 2022) transforms the recommendation task into a natural language processing task, generating personalized recommendations through prompts. VIP5 (Geng et al., 2023) leverages pre-trained multimodal models to capture complementary information from visual and textual data, providing more personalized and relevant recommendations to users. FedPA (Zhang et al., 2024c) and TransFR (Zhang et al., 2024b) integrate FM with FedRec. FedPA introduces a personalized adapter to fine-tune the FM, allowing each client to learn individual user preferences in a lightweight manner. In contrast, TransFR leverages FMs to generate universal textual representations, replacing traditional ID-based item representations and enhancing the system’s adaptability to new domains and its ability to handle cold-start issues.
3. Problem Formulation
Let $\mathcal{U}=\{u\}$ and =$\mathcal{I}=\{i\}$ be the set of users and items correspondingly, we assume that the users’ interaction history data is implicit feedback, represented by $\mathbf{R}=[\mathbf{r}_{1},\mathbf{r}_{2},\ldots,\mathbf{r}_{n}]^{T}\in\{0,1\}^{|\mathcal{U}|\times|\mathcal{I}|},$ where $r_{u,i}=1$ if user $u$ has interacted with item $i$, and $r_{u,i}=0$ otherwise. For user $u$, we use $\mathbf{r}_{u}\in\{0,1\}^{|\mathcal{I}|}$ to denote their interaction history and $\mathcal{I}_{u}$ to represent the set of items they have interacted with. Moreover, we define $\mathcal{O}=\{(u,i)|\text{user }u\text{ has interacted with item }i\}$ to index the items each user has interacted with. Each item in $\mathcal{I}$ has a set of modality features $\mathcal{M}$. The $m$-th modality feature is represented as $\mathbf{E}_{m}=[\mathbf{e}_{1,m},\mathbf{e}_{2,m},\dots,\mathbf{e}_{|\mathcal{I}|,m}]$. To capture user diversity, FedMR also introduces a set of modality fusion strategies $\mathcal{G}=\{g_{1},g_{2},\cdots,g_{|\mathcal{G}|}\}$. In this work, we focus on the visual and textual modalities, denoted as =$\mathcal{M}=\{v,c\}$, respectively. It is important to note that FedMR is not limited to these two modalities and the existing fusion strategies, it can include additional modalities or modify the fusion strategies as needed.
3.1. Objective of FedRec
In classic FedRec with implicit feedback, each client only has the user’s historical interaction data $\mathbf{r}_{u}$. The goal is to train a personalized recommendation model for each user while protecting user privacy, providing personalized recommendations by predicting possible scores $\hat{\mathbf{r}}_{u}$ for items the user has not yet interacted with, based on their observed historical data $\mathbf{r}_{u}$. To measure the difference between the prediction $\hat{\mathbf{r}}_{u}$ and the truth $\mathbf{r}_{u}$, we have the following:
(1) $\min\sum_{u\in\mathcal{U}}\alpha_{u}\sum_{(u,i)\in\mathcal{O}}\mathcal{L}_{\text{recon}}(\hat{r}_{u,i},r_{u,i}).$
Here, the function $\mathcal{L}_{\text{recon}}$ measures the reconstruction error between $\hat{\mathbf{r}}_{u}$ and $\mathbf{r}_{u}$, and $\alpha_{u}$ represents the loss weight for the client $u$ ($u\in\mathcal{U}$), which is used to balance the client’s contribution to the overall training, and satisfies the condition $\sum_{u\in\mathcal{U}}\alpha_{u}=1$. By optimizing Eq. (1), FedRec aims to make the predicted item scores $\hat{\mathbf{r}}_{u}$ for each user $u$ as close as possible to their true interactions ${\mathbf{r}}_{u}$.
Figure 1. The framework of FedMR. FedMR utilizes a FM to process multimodal data, including visual and textual information, to enhance personalized recommendations. On the server side, FedMR encodes visual and textual data into fixed embeddings. On each client, a Mixing Feature Fusion Module dynamically combines these embeddings with ID embeddings, adjusting the fusion weights based on user interactions to generate personalized item embeddings, capturing fine-grained user diversity.
3.2. ID-based FedRec
Currently, most FedRec methods (Lin et al., 2020a, b; Liang et al., 2021; Perifanis and Efraimidis, 2022; Zhang et al., 2023c, d; Li et al., 2023a, 2024b) construct a global item embedding $\mathbf{D}$ to capture the items’ features and their relationships with neighboring items, forming $\hat{\mathbf{r}}_{u}$ as follows:
(2) $\hat{\mathbf{r}}_{u}=f(\mathbf{D};\theta_{u}),$
where the function $f$, driven by the user-specific parameters $\theta_{u}$, generates personalized predictions $\hat{\mathbf{r}}_{u}$ upon receiving $\mathbf{D}$. However, these methods overlook the rich item multimodal information. Therefore, we consider such methods as ID-based FedRec in this work.
4. Methodology
The aim of our proposed method FedMR is to integrate the multimodal features of items with ID embeddings within a federated environment, and incorporates these features into existing ID-based FedRec methods. By doing so, we transform current FedRec methods into FedMMR methods, allowing them to fully utilize the multimodal features of items to predict items that users have not yet viewed but may find interesting, while protecting user privacy.
4.1. Encoding Multimodal Information with FM
In RecSys, users typically interact with only a small subset of all available items, which are usually stored on the server. Therefore, in multimodal scenarios, we assume the server holds the complete modal information of the item set. FedMR begins by introducing a foundation model $h$ on the server side to encode the items’ modal information, producing feature embeddings for each modality:
(3) $\mathbf{V}=h(\mathbf{E}_{v}),\quad\mathbf{C}=h(\mathbf{E}_{c}).$
Eq. 3 has two main advantages: 1) It transforms the visual or textual modality $\mathbf{E}_{m}$ into a format that the recommendation model can process; 2) Compared to transmitting raw modality, the encoded feature embeddings require less communication overhead, thus improving communication efficiency. Moreover, FedMR also builds an id embedding $\mathbf{D}$ to capture the structure information of items.
4.2. Mixing Feature Fusion Module
Considering the diverse interests and behavior patterns of users, FedMR incorporates a Mixing Feature Fusion Module (MFFM) that adopts multiple modality fusion strategies $\{g_{j}\}$ from $\mathcal{G}$ to fuse the multimodal feature embeddings $\mathbf{V}\text{, }\mathbf{C}$, and the id embedding $\mathbf{D}$ :
(4) $\mathbf{F}_{j}=g_{j}(\mathbf{V},\mathbf{C},\mathbf{D};\gamma_{j}),$
where $\gamma_{j}$ is parameters of the fusion strategy $g_{j}$, and $\mathbf{F}_{j}$ is the feature embedding obtained after applying the $j$-th fusion strategy. To capture user diversity by enabling each strategy to focus on specific user groups based on their interaction histories $\mathbf{r}_{u}$, FedMR introduces a user-specific router $\phi_{u}$, driven by the parameters $\varphi_{u}$:
(5) $\mathbf{w}_{u}=\phi_{u}(\{\mathbf{F}_{j}\}_{j=1}^{|\mathcal{G}|};\mathbf{r}_{u},\varphi_{u}).$
The router $\mathbf{r}_{u}$ dynamically generates a weight vector $\mathbf{w}_{u}\in\mathbb{R}^{|\mathcal{G}|}$ based on the feature embeddings of items the user has interacted with, obtained through different fusion strategies, and assigns weights to each fusion strategy, i.e., the weight of the strategy $g_{j}$ is $w_{u,j}$. Then, we have the final fused features $\bar{\mathbf{F}}$ as follows:
(6) $\bar{\mathbf{F}}=\sum_{j=1}^{|\mathcal{G}|}w_{u,j}\mathbf{F}_{j}.$
According to Eq. (6), FedMR dynamically integrates and weights each user’s $u$ multimodal information and ID features to generate the personalized item embeddings $\bar{\mathbf{F}}$, enhancing the system’s ability to capture complex user interaction patterns.
4.3. Objective of FedMR
By replacing $\mathbf{D}$ in Eq. (2) with $\bar{\mathbf{F}}$ obtained from Eq. (6) and then substituting into Eq. (1), we have the objective of FedMR as follows:
(7) $\min_{\{\theta_{u}\},\{\varphi_{u}\},\{\gamma_{j}\},\mathbf{D}}\sum_{u\in\mathcal{U}}\alpha_{u}\sum_{(u,i)\in\mathcal{O}}\mathcal{L}_{\text{recon}}(f(\bar{\mathbf{F}};\theta_{u}),r_{u,i}).$
Eq. (7) defines the objective of FedMR and is compatible with most existing ID-based FedRec methods. By replacing the original reconstruction error calculation function and the values of $\{\alpha_{u}\}$ in these methods with Eq. (7), they can integrate the multimodal information of items into personalized recommendations while protecting user privacy. This leads to a more comprehensive understanding of item features and enhances the model’s grasp of user preferences.
4.4. Framework
Fig. 1 illustrates the overall framework of FedMR, which leverages a FM to process multimodal information for enhancing personalized recommendations. On the server side, FedMR uses item cover images as visual modality data and titles as textual modality data. These inputs are processed by the FM to generate fixed visual embeddings $\mathbf{V}$ and text embeddings $\mathbf{C}$. Since both the FM and the input data are fixed, these item feature embeddings do not need to be updated during training. Additionally, the server maintains an ID embedding $\mathbf{D}$ to capture the structural information of the items.
Upon receiving the feature data $\mathbf{V}$, $\mathbf{C}$ and $\mathbf{D}$ from the server, the client filters the embeddings of items that the user has interacted with and performs fusion using the MFFM. Within this module, a learnable user-specific router $\phi_{u}$ dynamically assigns weights to the fused embeddings for further weighted fusion. The fusion strategies include a Sum strategy, a MLP (Taud and Mas, 2018) strategy, and a Gate strategy. The fused features $\bar{\mathbf{F}}$ are then fed into the prediction function $f$ to generate predicted scores $\hat{\mathbf{r}}_{u}$ for items, and the local model parameters are optimized by minimizing reconstruction loss.
During the entire process, only the ID embedding $\mathbf{D}$ and the gradients of fusion strategies in the MFFM are sent back to the server for aggregation and updates, hereby protecting user privacy while integrating shared information across clients. The core of FedMR lies in dynamically adjusting weights in the MFFM based on the user’s local interactions $\mathbf{r}_{u}$, thereby enhancing the diversity and personalization of the RecSys while ensuring privacy protection. As a result, FedMR is compatible with existing ID-based FedRec methods, allowing multimodal item information to be integrated into these methods. Furthermore, FedMR can be extended by incorporating more modalities or adjusting fusion strategies as needed.
Algorithm 1 FedMR
Input: $\mathbf{R}$, $\eta$, $T$, $A$
Initialize: $\{\theta_{u}\}$, $\{\varphi_{u}\}$, $\{\gamma_{j}\}$, $\mathbf{D}$
GlobalProcedure:
1:  $\mathbf{V}\leftarrow h(\mathbf{E}_{v})$, $\mathbf{C}\leftarrow h(\mathbf{E}_{c})$ according to Eq. (3);
2:  for $t=1,2,\dots,T$ do
3:     $S_{t}$leftarrow randomly select $n_{s}$ from $n$ clients;
4:     for all client index $u\in S_{t}$ do
5:        $\hat{\mathbf{r}}_{u}$, $\mathbf{D}^{(u)}$, $\{\nabla^{(u)}_{j}\gamma\}$leftarrow ClientUpdate($\varphi$, $\theta$);
6:     end for
7:     $\mathbf{D}\leftarrow\sum_{u\in S_{t}}\alpha_{u}\mathbf{D}^{(u)}$;
8:     $\gamma_{j}\leftarrow\gamma_{j}-\eta\sum_{u\in S_{t}}\alpha_{u}\nabla^{(u)}_{j}\gamma$, $j=1,2,\cdots,|\mathcal{G}|$;
9:  end for
10:  return: $\hat{\mathbf{R}}=[\hat{r}_{1},\hat{r}_{2},\dots,\hat{r}_{n}]^{T}$
ClientUpdate:
1:  $\nabla^{(u)}_{j}{\gamma}\leftarrow 0$, $j=1,2,\cdots|\mathcal{G}|$;
2:  for $a=1,2,\dots,A$ do
3:     Compute partial gradients $\nabla_{\theta_{u}}$, $\nabla_{\varphi_{u}}$, $\nabla_{\mathbf{D}}$ and $\{\nabla^{(u)}_{\gamma_{j}}\}$ by differentiating Eq. (7);
4:     Update $\theta_{u}\leftarrow\theta_{u}-\eta\nabla_{\theta_{u}}$;
5:     Update $\varphi_{u}\leftarrow\varphi_{u}-\eta\nabla_{\varphi_{u}}$;
6:     Update $\mathbf{D}\leftarrow\mathbf{D}-\eta\nabla_{\mathbf{D}}$;
7:     Update $\gamma_{j}\leftarrow\gamma_{j}-\eta\nabla^{(u)}_{\gamma_{j}}$, $j=1,2,\cdots|\mathcal{G}|$;
8:     Accumulate $\nabla^{(u)}_{j}{\gamma}\leftarrow\nabla^{(u)}_{j}{\gamma}+\nabla^{(u)}_{\gamma_{j}}$, $j=1,2,\cdots|\mathcal{G}|$;
9:     Compute $\hat{\mathbf{r}}_{u}$ according to Eq. (2);
10:  end for
11:  return: $\hat{\mathbf{r}}_{u}$, $\mathbf{D}$, $\{\nabla^{(u)}_{j}{\gamma}\}$
4.5. Algorithm
To optimize the FedMR objective function defined in Eq. (7), we propose an alternating optimization algorithm for training FedMR. The detailed steps of the algorithm are presented in Alg. 1.
The FedMR framework consists of two main components: the GlobalProcedure and the ClientUpdate. In the GlobalProcedure, the server first uses a FM $h(\cdot)$ to encode the multimodal data, generating the corresponding embeddings $\mathbf{V}$ and $\mathbf{C}$. The server then randomly selects a subset of clients to participate in each training round to reduce the risk of privacy leakage caused by clients participating in consecutive rounds (Li et al., 2023a). For each selected client $u\in S_{t}$, the client update process is executed to obtain the local prediction result $\hat{\mathbf{r}}_{u}$, the local updated id embeddings $\mathbf{D}^{(u)}$, and the local accumulated gradient $\{\nabla^{(u)}_{j}\gamma\}$. The server aggregates the results returned by the clients in a weighted manner, and updates the global shared embedding $\mathbf{D}$ and fusion strategy parameters $\{\gamma_{j}\}$.
During the ClientUpdate, each client calculates the gradients of the model parameters, such as $\nabla_{\theta_{u}}$, $\nabla_{\varphi_{u}}$, $\nabla_{\mathbf{D}}$, and $\nabla^{(u)}_{\gamma_{j}}$, based on the FedMR objective function in Eq. (7) and updates these parameters using gradient descent (Amari, 1993). Additionally, each client accumulates the local gradient $\nabla^{(u)}_{\gamma_{j}}$ into $\nabla^{(u)}_{j}\gamma$ for the global update of the fusion strategy parameters. By incorporating a FM $h(\cdot)$ to handle multimodal data $\mathcal{M}$, FedMR achieves efficient personalized multimodal recommendation while preserving user privacy.
4.6. Complexity Analysis
Since FedMR is compatible with various FedRec methods, we focus on the time and space complexity introduced by incorporating multimodal data of items in this work. Assuming $n$ users, $m$ items and $q$ experts, FedMR , where the multimodal information of the items is encoded on the server side into feature embeddings $\mathbf{V}\in\mathbb{R}^{m\times d}$ and $\mathbf{C}\in\mathbb{R}^{m\times d}$ using a foundation model, with $d\ll\{n,m\}$. Since these embeddings remain fixed during training, they contribute a space complexity of $O(md)$ while having negligible impact on time complexity. FedMR also maintains an ID embedding $\mathbf{D}\in\mathbb{R}^{m\times d}$ to capture structural information of the items, with a space complexity of $O(md)$. During each communication round, the server aggregates and updates local data from the clients, requiring a time complexity of $O(\alpha_{u}md(q+1))$, while each client performs the Mixing Feature Fusion Module to fuse features, resulting in a total client-side space complexity of $O(md+K)$ and a time complexity of $O(P)$, where $K$ and $P$ represents the size and cost time of the component parameters, such as Router, MLP, Gate, etc, respectively. Therefore, FedMR requires a total space complexity of $O(n(md+K)+3md)$ and a total time complexity of $O(nmd(q+1)+nP)$.
Table 1. The statistical information of the datasets used in our work includes the following: #Interactions represents the number of observed interactions. #Users denotes the number of users. #Items indicates the number of items.
Dataset #Users #Items #Interactions Sparsity
KU 204 560 3488 96.95%
Dance 10231 1676 80086 99.53%
Food 5990 1125 36482 99.46%
Movie 15908 2528 111091 99.72%
4.7. Privacy Preservation
FedMR uses a distributed architecture based on FL to effectively reduce the risk of user privacy leakage. Except for the Mixing Feature Fusion Module, all other components of FedMR have the same privacy risk level as existing ID-based federated recommendation methods. In the Mixing Feature Fusion Module, only the locally updated ID embeddings $\mathbf{D}^{(u)}$ and the local accumulated gradients ${\nabla^{(u)}_{j}\gamma}$ are sent to the server, without revealing any raw user data, also maintaining the same communication privacy protection strategy as current methods (Li et al., 2023a; Zhang et al., 2023c; Li et al., 2024b; Feng et al., 2024; Wu et al., 2024). To protect user privacy, randomly selecting participants in each communication round and avoiding the same participants in consecutive rounds can effectively reduce the risk of privacy leakage (Chai et al., 2020; Li et al., 2023a). Additionally, FedMR can incorporate techniques such as differential privacy (Abadi et al., 2016) and stochastic gradient clipping to further enhance privacy protection. We also conduct an ablation study to evaluate the effectiveness of FedMR when additional privacy-preserving measures are applied.
5. Experiments
5.1. Datasets
In this study, we conducte comprehensive experiments on four downstream datasets from Ninerec [^1] The dataset publicly available on: https://github.com/westlake-repl/NineRec (Zhang et al., 2024a) to evaluate the performance and effectiveness of the proposed method FedMR: KU, Bili_Food (Food), Bili_Dance (Dance), and Bili_Movie (Movie). Table 1 provides the detailed statistics of these datasets. The KU dataset is derived from KuaiShou, a well-known short video platform in China, where each entry represents a short video. The Food, Dance, and Movie datasets are collected from the BiliBili platform, focusing on food and cooking videos, dance-related videos, and movie-related videos, respectively. Each dataset presents unique content types and user interaction patterns, but all share a common challenge of high sparsity, with sparsity levels exceeding 95%. For our experiments, we include only users who had interacted with at least 5 items in each dataset to avoid the cold-start problem. Additionally, for missing images in the datasets, we fill them with the average values of visual features; for missing titles, we set ”The title is missing.”
Table 2. Performance comparison of Centralized (Cent) and Federated (Fed) Methods on four real-world datasets using HR@50 and NDCG@50. The federated methods are evaluated in both their original versions (vanilla) and enhanced with FedMR (w/ FedMR). The percentages in parentheses indicate the improvement of the version with FedMR compared to the original version.
KU Dance Food Movie
Method HR@50 NDCG@50 HR@50 NDCG@50 HR@50 NDCG@50 HR@50 NDCG@50
VBPR 0.4412 0.1547 0.2306 0.0747 0.2472 0.0790 0.1429 0.0486
BM3 0.4167 0.1595 0.2654 0.0856 0.2563 0.0820 0.1917 0.0666
Cent MGCN 0.3382 0.1508 0.2363 0.0767 0.2409 0.0777 0.1697 0.0574
vanilla 0.1127 0.0562 0.0782 0.0201 0.0696 0.0191 0.0552 0.0158
FedAvg w/ FedMR 0.5147 (356.70% $\uparrow$) 0.1491 (165.30% $\uparrow$) 0.1007 (28.77% $\uparrow$) 0.0261 (29.85% $\uparrow$) 0.0948 (36.21% $\uparrow$) 0.0268 (40.31% $\uparrow$) 0.0564 (2.17% $\uparrow$) 0.0160 (1.27% $\uparrow$)
vanilla 0.2255 0.0643 0.0676 0.0177 0.0820 0.0219 0.0581 0.0163
FCF w/ FedMR 0.5539 (145.63% $\uparrow$) 0.1597 (148.37% $\uparrow$) 0.0699 (3.40% $\uparrow$) 0.0182 (2.75% $\uparrow$) 0.0918 (11.95% $\uparrow$) 0.0247 (12.79% $\uparrow$) 0.0667 (14.80% $\uparrow$) 0.0180 (10.44% $\uparrow$)
vanilla 0.1176 0.0249 0.0221 0.0055 0.0548 0.0163 0.0185 0.0054
FedNCF w/ FedMR 0.1961 (66.75% $\uparrow$) 0.0609 (144.58% $\uparrow$) 0.0708 (220.36% $\uparrow$) 0.0187 (240.00% $\uparrow$) 0.0843 (53.83% $\uparrow$) 0.0243 (49.08% $\uparrow$) 0.0406 (119.46% $\uparrow$) 0.0125 (131.48% $\uparrow$)
vanilla 0.0637 0.0230 0.0252 0.0068 0.0327 0.0093 0.0170 0.0042
PFedRec w/ FedMR 0.2157 (238.62% $\uparrow$) 0.0616 (167.83% $\uparrow$) 0.0392 (55.56% $\uparrow$) 0.0092 (35.29% $\uparrow$) 0.0836 (155.66% $\uparrow$) 0.0207 (122.58% $\uparrow$) 0.0413 (142.94% $\uparrow$) 0.0132 (214.29% $\uparrow$)
vanilla 0.4706 0.1086 0.0682 0.0186 0.0696 0.0191 0.0453 0.0136
Fed FedRAP w/ FedMR 0.4824 (2.51% $\uparrow$) 0.1305 (20.17% $\uparrow$) 0.0895 (31.19% $\uparrow$) 0.0247 (32.91% $\uparrow$) 0.1142 (64.08% $\uparrow$) 0.0294 (53.93% $\uparrow$) 0.0586 (29.36% $\uparrow$) 0.0199 (46.32% $\uparrow$)
5.2. Baselines
To comprehensively evaluate the performance of FedMR, we select 3 cutting-edge centralized MMRec methods, and 5 ID-based federated methods commonly used in FedRec tasks with implicit feedback to serve as baselines in our study:
VBPR (He and McAuley, 2016) incorporates visual features into personalized ranking to improve recommendation accuracy based on implicit feedback.
BM3 (Zhou et al., 2023a) effectively integrates multimodal features through a self-supervised learning framework.
MGCN (Yu et al., 2023) utilizes a multi-view graph convolutional network to separately capture the multimodal features of users and items.
FedAvg (McMahan et al., 2017) is the earliest federated learning method, performing weighted aggregation based on the relative sizes of local client data.
FCF (Ammad-Ud-Din et al., 2019) introduces collaborative filtering into federated learning, assigning equal weights to all clients for average aggregation.
FedNCF (Perifanis and Efraimidis, 2022) updates user embeddings locally on each client and synchronizes item embeddings on the server for global updates.
PFedRec (Zhang et al., 2023c) aggregates item embeddings from clients on the server side and uses them during evaluation to achieve bilateral personalized recommendations.
FedRAP (Li et al., 2023a) achieves fine-grained personalized modeling of users and items through additive personalization.
The methods VBPR, BM3, and MGCN are three centralized benchmark MMRec methods that demonstrate baseline performances of the used datasets in a centralized recommendation setting. The methods FedAvg, FCF, FedNCF, PFedRec, and FedRAP are five ID-based FedRec methods. We conducted experiments on both the vanilla versions of these ID-based FedRec methods and their multimodal versions integrated with our proposed method FedMR to demonstrate the effectiveness of FedMR. Implementations of all used baseline methods are publicly available through their papers. All methods are implemented using PyTorch (Paszke et al., 2019), and all experiments are run on a machine with a 3.60 GHz Intel Xeon Gold 6244 CPU, 3 Quadro RTX 6000 GPUs, and 395 GB of RAM.
5.3. Experimental Setting
Following previous studies (Zhang et al., 2023c; Li et al., 2023a), we perform negative sampling for each positive sample in the training set and use the leave-one-out strategy for validation. We conduct hyperparameter tuning for all baseline methods on each dataset used, selecting the learning rate $\eta$ and adjusting the hyperparameters in FedRAP’s objective function from $\{10^{i}\mid i=-4,\ldots,-1\}$. To ensure fairness, we use the CLIP (Radford et al., 2021a) implemented by the Open CLIP (Ilharco et al., 2021; Radford et al., 2021b; Schuhmann et al., 2022) as the foundation model in our experiments, loading the ViT-B-32 parameters trained by OpenAI to process multimodal information. It should be noted that the foundation model in FedMR can be replaced with other models as needed. Additionally, we add a feature mapping layer before the input fusion layer of the multimodal embeddings and ID embeddings, mapping the feature dimensions output by the foundation model to a fixed latent embedding dimension, i.e. 64, to improve storage efficiency. We fix the training batch size at 2048 for all methods, and for methods with hidden layers, the number of layers is set to 3. For each federated baseline, we adopt the same aggregation strategy to update the global part of FedMR. In the main experiments, all clients participated in the training of both the original versions and the FedMR versions of all federated baselines, and no additional privacy protection measures are applied. Additionally, for these federated baselines, the number of local training epochs was set to 10. Except for loading the foundation model, we do not use any pre-training strategies for any methods.
5.4. Evaluation
In our experiments, we treat the observed interactions as positive samples and the rest as negative samples. We sort the predicted interaction scores between users and all candidate items in descending order, masking the observed positive samples from the training set. Then, we evaluate the recommendation performance using two widely-used metrics: Hit Rate (HR@K) and Normalized Discounted Cumulative Gain (NDCG@K), which are formally defined in (He et al., 2020). In the experiments, we set $K=50$.
5.5. Comparison Analysis
Table 2 shows the performance of both centralized and federated methods across the four datasets. The results highlight that FedMR significantly boosts the performance of the used FedRec methods, particularly in the KU dataset, where HR@50 for FedAvg, FCF, PFedRec, and FedRAP exhibits notable improvements. This indicates that FedMR effectively enhances traditional ID-based FedRec methods by integrating multimodal item information. The performance improvements in other datasets further demonstrate FedMR’s ability to alleviate data sparsity by better leveraging users’ interaction histories alongside multimodal data for more personalized recommendations, while protecting user privacy. Compared to centralized MMRec methods, FedRec approaches using FedMR achieve similar or even superior performance while maintaining privacy, e.g., in the KU dataset, FedAvg with FedMR surpasses all centralized methods in terms of HR@50. Federated methods with FedMR also perform strongly across other datasets, reinforcing FedMR’s capability to improve recommendation quality without sacrificing user privacy.
Table 2 further shows that by incorporating multimodal information, FedMR provides a more comprehensive representation of item features. Compared to ID-based FedRec, the use of multimodal data enables more accurate predictions of user preferences. Additionally, the MFFM of FedMR dynamically adjusts the weights of visual, textual, and ID features based on each user’s dependence on these modalities. This flexible approach to feature fusion avoids the drawbacks of single fusion strategy methods, allowing the model to better capture complex relationships between modalities and further enhance personalized recommendation accuracy.
5.6. Ablation Study
To validate the effectiveness of each component in FedMR, we conducted experiments using several variants of the system. Specifically, we independently applied the Sum, MLP, and Gate strategies for the fusion of multimodal features and ID features, to assess the impact of dynamic weighted fusion based on users’ interaction history $\mathbf{r}_{u}$, $\forall u\in\mathcal{U}$. Additionally, we introduce random Gaussian noise into the gradients uploaded to the server to evaluate the potential performance impact of this privacy-preserving measure. We also examine the distribution of multimodal and ID features within the same feature space. These ablation studies demonstrate the effectiveness of each component of FedMR.
Figure 2. Visualization of multimodal features and ID features mapped into a shared representation space.
5.6.1. Ablation Study on Modality Distributions
To investigate the distribution relationship between multimodal features and ID features, we select FedRAP (Li et al., 2023a) as the backbone for FedMR in this study. FedRAP maintains both globally shared and user-specific ID features, which helps demonstrate how different ID features affect the distribution within the feature space.
Fig. 2 visualizes the distribution of modality features and ID features after training FedRAP with FedMR, on the KU dataset. The first subfigure in Fig. 2 shows that when the multimodal features of items and globally shared ID features are mapped into a shared representation space, they remain in distinct subregions. This phenomenon, known as the ”Modality Gap” (Liang et al., 2022; Zhang et al., 2023a), helps the model distinguish between different modalities, reducing noise interference and improving the robustness of multimodal data processing. The following two subfigures incorporate the personalized ID embeddings of the users 00 and $42$, respectively, highlighting the significant impact of personalized ID features on the distribution in the shared space. By comparing the first subfigure with the subsequent two, we observe that the introduction of personalized ID features causes the ID features to gradually expand towards the regions of visual and textual features, shrinking the modality gap. This facilitates the fusion of multimodal information and enhances the accuracy of personalized recommendations. Furthermore, by comparing the differences in modality gap changes for each user, it becomes evident that the model can adjust the feature distribution based on user-specific behaviors and preferences, thus better satisfying individual user needs.
This study underscores the importance of personalized ID features in federated recommendation systems. Due to privacy constraints, the model cannot access raw user data directly, but through the use of personalized ID features, the model can provide personalized recommendation services while maintaining user privacy.
Table 3. Performance comparison of FedRec Methods on HR@50 and NDCG@50 with and without multimodal and ID features on the KU dataset.
Method Metric FedMR w/o $\mathbf{V}$ w/o $\mathbf{C}$ w/o $\mathbf{D}$
FedAvg HR@50 0.5147 0.4069 (20.94% $\downarrow$) 0.3725 (27.63% $\downarrow$) 0.5094 (1.03% $\downarrow$)
NDCG@50 0.1491 0.1391 (6.71% $\downarrow$) 0.1287 (13.68% $\downarrow$) 0.1407 (5.66% $\downarrow$)
FCF HR@50 0.5539 0.1569 (71.67% $\downarrow$) 0.1569 (71.67% $\downarrow$) 0.1275 (76.99% $\downarrow$)
NDCG@50 0.1597 0.0486 (69.57% $\downarrow$) 0.0476 (70.19% $\downarrow$) 0.0287 (82.04% $\downarrow$)
FedNCF HR@50 0.1961 0.1569 (19.99% $\downarrow$) 0.1863 (5.00% $\downarrow$) 0.1602 (18.31% $\downarrow$)
NDCG@50 0.0609 0.0431 (29.23% $\downarrow$) 0.0541 (11.17% $\downarrow$) 0.0538 (11.66% $\downarrow$)
PFedRec HR@50 0.2157 0.2010 (6.82% $\downarrow$) 0.1225 (43.21% $\downarrow$) 0.0490 (77.27% $\downarrow$)
NDCG@50 0.0616 0.0557 (9.58% $\downarrow$) 0.0341 (44.64% $\downarrow$) 0.0191 (68.97% $\downarrow$)
FedRAP HR@50 0.3824 0.2647 (30.78% $\downarrow$) 0.2647 (30.78% $\downarrow$) 0.1863 (51.29% $\downarrow$)
NDCG@50 0.1305 0.0762 (41.61% $\downarrow$) 0.0734 (43.75% $\downarrow$) 0.0490 (62.43% $\downarrow$)
5.6.2. Ablation Study on Modality Efficiency
Table 3 presents a performance comparison of the five used ID-based FedRec methods with the inclusion of multimodal information. It evaluates these methods under the full model (FedMR) and with the removal of visual modality (w/o V), textual modality (w/o C), and ID features (w/o D), focusing on HR@50 and NDCG@50 metrics.
Firstly, removing the visual modality results in a noticeable performance drop across all methods, particularly for FCF (HR@50 decreased by 71.67%) and FedRAP (HR@50 decreased by 30.78%). This highlights the critical role of visual information in RecSys, especially for FCF. Visual features provide rich, intuitive image-based information that helps models better capture user preferences regarding the appearance of items. Without this visual context, models struggle to accurately infer user preferences, leading to a significant decline in recommendation performance. Secondly, removing the ID feature also causes a substantial decline in performance. The ID feature is essential for personalized recommendations, as it captures users’ historical interactions and reflects their past behaviors and interests. When the ID feature is removed, the model loses its primary mechanism for personalization, resulting in decreased accuracy and personalization of recommendations. This impact is especially pronounced in methods like PFedRec (HR@50 decreased by 82.82%) and FedRAP (HR@50 decreased by 51.92%), which heavily rely on ID features. In contrast, removing the text modality has a relatively smaller impact, though it still affects some methods. For instance, in FedAvg and PFedRec, removing the text modality led to HR@50 decreases of 27.63% and 43.21%, respectively. This indicates that text information plays an important role in capturing semantic relationships between user behaviors and item descriptions, helping the model better understand user needs and item characteristics, thereby enhancing recommendation performance.
In summary, all methods experienced performance degradation to varying degrees when any modality was removed. Considering Fig. 2, visual, text, and ID features are often complementary. Removing any one of these disrupts the balance, making it difficult for the model to fully represent user preferences and item characteristics. Additionally, different methods depend on different modalities to varying extents. For instance, FCF and PFedRec are more dependent on visual information, whereas FedAvg is less reliant on ID features. These differences in dependency lead to varying levels of performance decline when specific modalities are removed. Further experimental results are provided in the appendix.
Figure 3. Performance comparison of different feature fusion strategies in various FedRec methods combined with FedMR on the KU dataset.
5.6.3. Ablation Study on Fusion Strategy
Fig. 3 presents the HR@50 and NDCG@50 evaluation results for five used FedRec methods using different feature fusion strategies with FedMR on the KU dataset. The results indicate that different fusion strategies significantly impact performance. Notably, the Mix strategy, which dynamically combines multiple fusion methods based on user interaction history, outperforms others, while the Sum strategy shows poor performance. This trend is consistent across all metrics.
The Sum strategy simply adds different modality features without considering their complex interactions and inter-relationships. Since each modality’s importance varies, simple addition can lead to information loss or imbalanced weighting, making it difficult to capture relationships effectively. Additionally, RecSys often face a long-tail distribution, where a few popular items dominate user interactions (Park and Tuzhilin, 2008; Yin et al., 2012). The simplicity of the Sum strategy may lead to overfitting on popular items, resulting in poor performance on less interacted items and affecting overall model performance.
In contrast, the MLP strategy uses fully connected networks to learn nonlinear modality combinations, which better captures complex interactions between item modalities and ID features. However, its effectiveness is limited by network depth and the number of parameters. On the other hand, the Gate strategy learns adaptive weights to dynamically filter the contribution of each modality, optimizing the importance of different features.
The Mix strategy integrates all fusion methods and adjusts their weights dynamically based on user interaction history, allowing the model to capture more fine-grained personalized item representations. This approach enables more accurate matching of user preferences and effectively captures complex relationships between multimodal features, leading to superior performance.
This study highlights the importance of personalized item embeddings in FedRec. In FedMR, each client dynamically adjusts fusion strategy weights through the Mixing Feature Fusion Module, based on user interaction history, to generate personalized item embeddings, enhancing personalization and accuracy while maintaining user privacy, thus improving overall system performance.
Table 4. Performance comparison of FedRec methods combined with FedMR, with and without Gaussian noise (mean=0, variance=0.1) on KU dataset for privacy protection.
Method Metric FedMR w/ noise Degrade
FedAvg HR@50 0.5147 0.4969 3.46% $\downarrow$
NDCG@50 0.1491 0.1416 5.03% $\downarrow$
FCF HR@50 0.5539 0.5378 2.91% $\downarrow$
NDCG@50 0.1597 0.1402 12.21% $\downarrow$
FedNCF HR@50 0.1961 0.1680 14.33% $\downarrow$
NDCG@50 0.0609 0.0420 31.03% $\downarrow$
PFedRec HR@50 0.2157 0.1965 8.90% $\downarrow$
NDCG@50 0.0616 0.0586 4.87% $\downarrow$
FedRAP HR@50 0.3824 0.3725 2.59% $\downarrow$
NDCG@50 0.1305 0.1236 5.29% $\downarrow$
5.6.4. Privacy Protection
Table 4 presents the experimental results of the used FedRec methods combined with FedMR, comparing the performance with and without adding Gaussian noise (mean=0, variance=0.1) to the gradients of the global components uploaded to the server, to protect user privacy. The experiments are conducted on the KU dataset during feature fusion. The results indicate that all methods experienced some performance degradation after adding noise, particularly on the NDCG metric, which is more sensitive to ranking accuracy. This is because noise disturbs the gradient precision, which affects the model’s ability to update effectively.
Among the methods, FedNCF is the most sensitive to noise, with HR@50 dropping by 14.33% and NDCG@50 by 31.03%. As a complex neural network-based model, FedNCF depends on numerous parameters to capture nonlinear features. The introduction of noise disrupts these feature interactions, resulting in a significant decline in performance. In contrast, FedAvg and FedRAP show greater robustness to noise, with only minimal performance degradation. This suggests that these methods are more effective at maintaining recommendation accuracy while ensuring privacy protection.
In conclusion, although adding noise helps protect user privacy, its impact varies based on model complexity and feature interaction mechanisms. Simpler models like FedAvg and adaptive models like FedRAP are more resilient to noise, while complex neural network models like FedNCF experience more significant performance drops due to the noise interference.
6. Conclusion
This paper presents FedMR, a federated multimodal recommendation system that incorporates a mixing feature fusion module on each client. This module dynamically adjusts fusion weights based on user interaction history, generating personalized item embeddings that capture fine-grained user preferences, thereby enhancing personalized recommendations. FedMR allows flexible modification of fusion strategies and supports different modalities of item data as needed. Experimental results show that FedMR performs well on four multimodal recommendation datasets and is compatible with existing ID-based methods without requiring changes to the underlying framework. This study offers new directions for the development of federated recommendation systems, particularly in the application of multimodal data and foundation models.
Appendix A Algorithm Optimization
A.1. Derivation of the Update for $\gamma_{j}^{(u)}$
The update process for $\{\gamma_{j}^{(u)}\}_{j=1}^{|\mathcal{G}|}$ in Alg. 1 is based on gradient descent. The local update at each client $u$ during the iteration $t$ follows the following rule:
(8) $\gamma_{j}^{(t)}=\gamma_{j}^{(t-1)}-\eta\nabla_{j}^{(u)}\gamma^{(t)},$
where $\eta$ is the learning rate, and $\nabla_{j}^{(u)}\gamma^{(t)}$ represents the gradient of $\gamma_{j}^{(u)}$ at the iteration $t$ on client $u$.
To capture how $\gamma_{j}^{(u)}$ evolves over time, we can unroll this equation over several iterations. By recursively applying the update rule in Eq. (8), we obtain the following:
(9) $\gamma_{j}^{(t)}=\gamma_{j}^{(0)}-\eta\sum_{i=1}^{t}\nabla_{j}^{(u)}\gamma^{(i)},$
where $\gamma_{j}^{(0)}$ is the initial value of $\gamma_{j}^{(u)}$, and the second term is the sum of all gradients computed up to the $t$-th iteration. This expression shows that the updated value of $\gamma_{j}^{(u)}$ depends on its initial value and the accumulated gradient over $t$ iterations.
A.2. Accumulated Gradient
The accumulated gradient $\nabla_{j}^{(u)}\gamma$ at the client $u$ is the sum of the individual gradients across the iterations, which be expressed as:
(10) $\nabla_{j}^{(u)}\gamma=\frac{1}{\eta}\left(\gamma_{j}^{(0)}-\gamma_{j}^{(t)}\right).$
Eq. (10) represents the total change in $\gamma_{j}^{(u)}$ during the local training process. It relates the initial value $\gamma_{j}^{(0)}$ to the current value $\gamma_{j}^{(t)}$, scaled by the learning rate $\eta$.
A.3. Global Aggregation
In federated learning, the global model is updated by aggregating the local updates from multiple clients. For the parameter $\gamma_{j}$, this is achieved by summing the accumulated gradients $\nabla_{j}^{(u)}\gamma$ from all selected clients. The global update for $\gamma_{j}$ can be written as:
(11) $\gamma_{j}=\gamma_{j}-\eta\sum_{u\in\mathcal{S}_{t}}\alpha_{u}\nabla_{j}^{(u)}\gamma,$
where $\mathcal{S}_{t}$ is the set of clients selected during round $t$. This step updates the global parameter $\gamma_{j}$ by aggregating the contributions from each client’s local updates. The derived equations demonstrate that the update process is consistent with the principles of gradient-based optimization. This ensures that $\gamma_{j}^{(u)}$ is updated in an efficient and mathematically sound manner within the FL framework.
Table 5. Changes of parameter numbers in FedRec with and without FedMR on the KU Dataset, Compared to Centralized Recommendation Systems.
Method KU Dance Food Movie
Cent VBPR 127552 1482432 904320 2263616
BM3 1752576 8600576 5582848 12815872
MGCN 708864 2564800 1693888 3855104
Fed FedAvg Vanilla 35905 107329 72065 161857
FedMR 455620 1027012 744900 1463236
FCF Vanilla 35905 107329 72065 161857
FedMR 455620 1027012 744900 1463236
FedNCF Vanilla 108737 1535041 921665 1182689
FedMR 564292 2561988 1666500 3833924
PFedRec Vanilla 35905 107329 72065 161857
FedMR 455620 1027012 744900 1463236
FedRAP Vanilla 71745 214593 144065 323649
FedMR 742340 1885124 1320900 2757572
Appendix B Parameter Analysis
Table 5 illustrates the change in the number of learnable parameters on a single client in the KU dataset when FedMR is integrated into FedRec or not, compared to centralized recommendation systems. As shown in the table, the introduction of FedMR significantly increases the parameter size. This is because FedMR incorporates multimodal information (e.g., visual and textual features) alongside ID embeddings, increasing the model’s complexity and consequently requiring more parameters to learn.
Despite the increase in the number of learnable parameters, Table 2 demonstrates that FedMR significantly improves the recommendation performance of FedRec on the KU dataset. By combining multimodal features with ID embeddings, FedMR better leverages users’ interaction histories to capture the fine-grained personality of users. Additionally, the mixing feature fusion module in FedMR adjusts the weights of multimodal features based on users’ interaction history, greatly enhancing both the accuracy and personalization of recommendations. Moreover, compared to traditional FedRec methods, which rely solely on the ID embeddings, FedMR enables personalized modeling by integrating multimodal information, thereby further improving recommendation quality. While the model’s complexity increases, this trade-off results in significant performance gains.
As privacy protection is crucial in the federated settings, FedMR processes multimodal data locally on the client and only uploads essential parameters, e.g., ID embeddings and gradients, to the server, avoiding the exposure of raw user data. Although this adds computational overhead and increases parameter size on the client, these sacrifices are necessary to ensure privacy while also boosting recommendation performance. Thus, the increased learnable parameter size is not only a necessary cost for improved recommendation performance but also a vital measure for ensuring privacy.
Appendix C More Experiments
Figure 4. Performance comparsions of different FedRec methods with FedMR under varying client sampling ratios on HR@50 and NDCG@50.
C.1. Ablation Study on Client Sampling
To investigate the impact of introducing FedMR for multimodal data fusion on different FedRec methods under varying client sampling ratios, we conduct experiments on the five FedRec methods using the KU dataset. Fig. 4 presents the results for HR@50 and NDCG@50 as the client sampling ratio changes. Overall, as the client sampling ratio increases, most methods show improvements in HR@50 and NDCG@50, indicating that more client participation helps enhance the personalization of recommendations. However, the performance improvement varies across different methods. For PFedRec, the increase in HR@50 and NDCG@50 after introducing FedMR is relatively small, possibly due to its lower sensitivity to the number of participating clients. This suggests that PFedRec may rely more on the quality of individual client data rather than the number of clients. As a result, the advantages of FedMR’s multimodal fusion are not fully realized when the client count is low. Additionally, when the client sampling ratio is low, the model’s performance becomes unstable, indicating that insufficient data affects the stability of feature learning. This is especially true in multimodal scenarios, where a lack of sufficient modality information weakens the fusion effect, reducing recommendation accuracy.
Appendix D Limitations and Future Work
While FedMR has shown promising improvements in federated multimodal recommendation systems, several limitations remain:
  1. (1)
    Data Heterogeneity: Considering the federated settings, the distribution of user data varies significantly. Although FedMR mitigates this issue by integrating multimodal information, its performance is still limited when handling highly heterogeneous data. Future work could focus on introducing more robust personalization models or adaptive aggregation strategies to better address the differences between users.
  2. (2)
    Communication Overhead: Although FedMR reduces some communication costs by encoding multimodal data, processing multimodal features still results in high computational and communication costs. Future research could explore optimizing communication protocols or using model compression techniques to further minimize these costs.
  3. (3)
    Limited Flexibility in Multimodal Fusion: While FedMR supports dynamic adjustments to its fusion strategies, its current approach may lack flexibility when dealing with complex user behaviors. Future work could explore more adaptive and layered fusion methods to better capture the diverse preferences of users.
In conclusion, future developments could enhance the overall performance of federated recommendation systems by improving the handling of data heterogeneity, reducing communication costs, and optimizing multimodal fusion strategies.