### v3

# Dynamic Fusion Strategies for Federated Multimodal Recommendations

Abstract.
Delivering deeply personalized recommendations necessitates understanding user interactions with diverse multimedia features, but achieving this within the constraints of Federated Recommendation Systems (FedRec) is severely hampered by communication bottlenecks, user heterogeneity, and the complexity of privacy-preserving multimodal fusion. To this end, we propose FedMR, a novel multimodal FedRec framework centered around the Mixing Feature Fusion Module (MFFM). FedMR employs a two-stage process: (1) Server-side centralized multimedia content processing provides rich, shared item context using pre-trained models, mitigating limitations from client sparsity and resource constraints efficiently. (2) Client-Side Personalized Refinement, where the MFFM dynamically adapts these server-provided multimodal representations based on client-specific interaction patterns, effectively tailoring recommendations and resolving heterogeneity in user preferences towards different modalities. Extensive experiments validate that FedMR seamlessly enhances existing ID-based FedRecs, effectively transforming them into high-performing federated multimodal systems.
Federated Recommendation, Multimodal Learning, Personalization
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Collaborative filteringccs: Information systems Personalizationccs: Information systems Combination, fusion and federated search
1. Introduction
Recommendation Systems (RecSys) are integral components of modern digital platforms, guiding users through vast item catalogs by suggesting relevant content (Li and Long, 2024). Traditional RecSyss rely heavily on collaborative filtering signals derived from user-item interaction histories (Chen et al., 2017). However, the rich information embedded within item multimedia content, such as product images, video trailers, textual descriptions, and audio samples, offers a powerful complementary signal (Deldjoo et al., 2020; Zhou et al., 2023b). Leveraging these multimodal cues allows Multimodal Recommendation Systems (MMRec) (He and McAuley, 2016; Wei et al., 2019; Zhou et al., 2023a; Yu et al., 2023) to gain a deeper understanding of item characteristics and user preferences, capture subtle nuances missed by interaction data alone, significantly alleviate the cold-start problem for new items (Pan et al., 2022), and ultimately enhance recommendation accuracy and user experience.
Despite their effectiveness, centralized MMRec approaches typically require aggregating vast amounts of user interaction data and potentially sensitive item features on central servers (Liu et al., 2024). This practice raises significant privacy concerns, particularly with the enforcement of stringent data protection regulations like the GDPR (Voigt and Von dem Bussche, 2017). To address these concerns, Federated Learning (FL) (McMahan et al., 2017) has emerged as a compelling paradigm, enabling collaborative model training across distributed clients, e.g., user devices, without exposing raw local data. Federated Recommendation Systems (FedRec) (Yang et al., 2020; Wang et al., 2022) adapt FL principles to the recommendation domain, offering a promising path towards privacy-preserving personalization.
However, integrating rich multimedia information into the FL setting poses distinct challenges beyond traditional ID-based FedRecs (Feng et al., 2024). First, processing raw, high-dimensional multimedia data on resource-constrained client devices is often computationally prohibitive, hindering feature extraction and user participation (Li et al., 2024a). Second, transmitting large multimedia features incurs significant communication overhead and raises privacy concerns, as shared features might reveal sensitive information (Chai et al., 2020; Li et al., 2023a). Third, data heterogeneity (Yuan et al., 2024; Li et al., 2019) extends to user preferences for different modalities, such as visual vs. textual (Liu et al., 2019; Niu et al., 2021). This makes a single, global fusion strategy suboptimal for personalized multimodal understanding (Feng et al., 2024; Li et al., 2024a). Finally, client-side data sparsity (Sun et al., 2024) limits item understanding, especially for cold-start items (Zhang et al., 2024c). While multimedia data could alleviate this (He and McAuley, 2016; Zhou et al., 2023a), its utilization is obstructed by the aforementioned constraints. Hence, designing a FedRec that integrate multimodal data while balancing efficiency, privacy, heterogeneity, and sparsity is thus a critical research problem (Feng et al., 2024; Li et al., 2024a).
To address these interconnected challenges, we propose FedMR, a novel Federated Multimodal Recommendation framework. FedMR utilizes a collaborative server-client architecture where the server leverages powerful pre-trained models for centralized, efficient encoding of item modalities, creating rich, shared contextual representations that combat client-side item understanding limitations due to sparsity. Crucially, FedMR introduces the Mixing Feature Fusion Module (MFFM), operating on the client-side for dynamic and personalized adaptation. The MFFM empowers each client to learn user-specific weights for combining different fusion perspectives of the server-provided multimodal representations and collaborative signals from ID embeddings, guided by the client’s unique interaction history. This adaptive multimodal fusion directly tackles data heterogeneity by tailoring the final item representation to individual user preferences and local data characteristics, leading to more accurate and personalized recommendations while preserving privacy.
Our main contributions are summarized as follows:
  • Framework Innovation: We propose FedMR, a novel federated architecture employing centralized server-side processing of item modalities to enrich item context and alleviate client-side limitations while protecting user privacy.
  • Dynamic Feature Fusion: We introduce the Mixing Feature Fusion Module for dynamic, personalized adaptation of multimodal representations, effectively addressing data heterogeneity and diverse modality preferences.
  • Seamless Compatibility and Enhancement: FedMR seamlessly integrates with existing ID-based FedRec frameworks, upgrading them with multimodal capabilities while requiring minimal modification and boosting performance.
  • Comprehensive Empirical Validation: Extensive experiments on real-world multimodal datasets validate FedMR’s effectiveness, showing significant improvements over FedRec baselines. We also provide code for reproducibility.
2. Related Work
2.1. Federated Recommendation Systems
FedRecs aim to provide personalized recommendations while ensuring user privacy and data security (Yang et al., 2020). Significant progress has been made in FedRec across various tasks, including rating prediction (Lin et al., 2020a, b; Liang et al., 2021; Lin et al., 2022; Zhang et al., 2023d; Li et al., 2024b), top-k recommendation (Ammad-Ud-Din et al., 2019; Perifanis and Efraimidis, 2022; Luo et al., 2022; Zhang et al., 2023b; Dong et al., 2022; Li et al., 2022a; Qu et al., 2023), and cold-start scenarios (Liu et al., 2021b; Meihan et al., 2022; Wahab et al., 2022; Zhang et al., 2024c, d; Li et al., [n. d.]). However, many existing FedRec approaches (Huang et al., 2020; Flanagan et al., 2021; Wu et al., 2022; Zhu et al., 2022; Imran et al., 2023; Yuan et al., 2023a, 2024) primarily use ID embeddings for item representations, thereby neglecting rich multimodal information like associated text and images. Recent efforts have started applying multimodal learning (Baltrušaitis et al., 2018) to FedRec, called Federated Multimodal Recommendation System (FedMMRec). AMMFRS (Feng et al., 2024) employs an attention mechanism (Niu et al., 2021) to weight and fuse multimodal data based on feature importance. In contrast, FedMMR (Li et al., 2024a) offloads heavy multimodal computation to the server and employs secure cross-client convolution with secret sharing to capture high-order interactions privately. Other works focus on enhancing personalization within FedRec. PFedRec (Zhang et al., 2023c) introduced dual personalization for both users and items. FedRAP (Li et al., 2023a) refined this with additive personalization. FedDAE (Li et al., 2024c) uses a variational autoencoder and a gating network to adaptively fuse shared and personalized item knowledge. Furthermore, recent works explore integrating Foundation Models (FMs) directly within FedRec. FedPA (Zhang et al., 2024c) introduces personalized adapters for lightweight, client-specific FM fine-tuning. TransFR (Zhang et al., 2024b) uses FMs to generate universal textual item representations, replacing traditional IDs to improve adaptability and address cold-start issues. Despite these advancements in personalization and FM integration, methods predominantly relying on ID embeddings or single-modality representations still limit the full utilization of rich, multi-faceted multimodal item information within the federated setting (Che et al., 2023).
2.2. Multimodal Recommendation Systems
Traditional Recommendation Systems (RecSys) primarily rely on user-item interactions (Yu et al., 2023). In contrast, Multimodal Recommendation systems (MMRec) enhance accuracy and user experience by leveraging multiple data modalities like text, images, audio, and video (Zhou et al., 2023b). Item IDs, often used as unique identifiers, can capture collaborative signals and structural relationships (Yuan et al., 2023b; Liu et al., 2023). In MMRec, these IDs are frequently combined with multimodal data as auxiliary information to create richer item representations (Wang et al., 2024; Zhang et al., 2024e). This integration helps capture a more comprehensive understanding of users’ interests and preferences, ultimately improving recommendation quality. A typical MMRec process involves two main stages: feature extraction and modality fusion (Guo et al., 2019; Gao et al., 2020). Feature extraction (Guo et al., 2019) aims to obtain representative features from various modalities. Foundation Models (FMs), typically pre-trained on large-scale unlabeled datasets (Bommasani et al., 2021), have become increasingly prominent for this task. They can be adapted for recommendation through various techniques, including representation transfer (Du et al., 2020; Liu et al., 2021a, 2022; Wei et al., 2023), direct fine-tuning (Zhang et al., 2021; Li et al., 2022b, 2023b; Sun et al., 2023; Wang et al., 2023), parameter-efficient adapter tuning (Hou et al., 2022; Bian et al., 2023; Geng et al., 2023; Fu et al., 2024), or prompt-based tuning (Shin et al., 2020; Li and Liang, 2021; Zhou et al., 2022; Zhang and Wang, 2023; Xu et al., 2024). Several studies leverage FMs to enhance the content understanding capabilities of MMRec, often using cross-modal alignment or generation techniques. CLIP (Radford et al., 2021a; Wu et al., 2024), for example, uses contrastive learning (Le-Khac et al., 2020) between language and images to obtain powerful multimodal representations. M6-Rec (Cui et al., 2022) is an FM-based system tailored for e-commerce recommendations. P5 (Geng et al., 2022) frames recommendation as a natural language task solvable via prompts. VIP5 (Geng et al., 2023) utilizes pre-trained multimodal models to capture complementary visual and textual information for personalized recommendations. Following feature extraction, modality fusion (Gao et al., 2020) focuses on effectively integrating these features to enhance overall performance. Recommendation models then combine these fused multimodal features with interaction data to predict user preferences (Tao et al., 2020; Wu et al., 2021; Zhou and Shen, 2023). However, realizing the full potential of MMRec within the constraints of FL, especially regarding personalized and adaptive fusion across heterogeneous clients, necessitates specialized frameworks are needed to provide personalized recommendation with the privacy protection.
3. Problem Formulation
Let $\mathcal{U}$ and $\mathcal{I}$ denote the sets of users and items, respectively. User-item interactions are represented by a binary matrix $\mathbf{R}\in\{0,1\}^{|\mathcal{U}|\times|\mathcal{I}|}$, where $r_{u,i}=1$ signifies an interaction between user $u\in\mathcal{U}$ and item $i\in\mathcal{I}$, and $r_{u,i}=0$ otherwise. The interaction history for user $u$ is denoted by the vector $\mathbf{r}_{u}\in\{0,1\}^{|\mathcal{I}|}$, and the set of observed interactions is $\mathcal{O}=\{(u,i)\mid r_{u,i}=1\}$. Each item $i\in\mathcal{I}$ is associated with features from a set of modalities $\mathcal{M}$. For the $m$-th modality, these features are represented by a matrix $\mathbf{E}_{m}\in\mathbb{R}^{|\mathcal{I}|\times d_{m}}$, where the vector $\mathbf{e}_{i,m}\in\mathbb{R}^{d_{m}}$ is the feature embedding for item $i$ in modality $m$. To handle multimodal information effectively, we introduce a set of modality fusion strategies $\mathcal{G}=\{g_{1},\dots,g_{|\mathcal{G}|}\}$ for dynamically combining features. We primarily focus on visual $v$ and textual $c$ modalities, i.e., =$\mathcal{M}=\{v,c\}$, although the FedMR framework is extensible to other modalities and fusion mechanisms.
4. Methodology
FedMR aims to seamlessly integrate multimodal item features with ID embeddings within a federated environment, enhancing existing ID-based FedRecs, to predict items that users have not yet explored but may find interesting. It centralizes multimodal data processing on the server for efficiency and rich item context, then uses dynamic client-side personalization via MFFM to handle heterogeneity.
4.1. Preliminaries
FedRec systems aim to collaboratively train recommendation models while preserving user privacy. The standard objective is to minimize the aggregated reconstruction loss across clients $u\in\mathcal{U}$:
(1) $\min\sum_{u\in\mathcal{U}}\alpha_{u}\sum_{(u,i)\in\mathcal{O}}\mathcal{L}_{\text{recon}}(\hat{r}_{u,i},r_{u,i}),$
where $\mathcal{L}_{\text{recon}}$ is the loss function, and $\alpha_{u}$ is the weight of client $u$’s contribution. Many conventional FedRec methods (Lin et al., 2020a, b; Liang et al., 2021; Perifanis and Efraimidis, 2022; Zhang et al., 2023d) rely solely on a global item ID embedding matrix $\mathbf{D}$ and user-specific parameters $\theta_{u}$ to predict a score $\hat{r}_{ui}$ using a prediction function $f$:
(2) $\hat{r}_{ui}=f(\mathbf{D}_{i};\theta_{u}).$
Here, $\mathbf{D}_{i}$ represents the ID embedding for item $i$. These methods, so-called ID-based FedRecs, primarily capture collaborative signals and structural relationships (Yuan et al., 2023b), but often overlook rich multimodal item content, which motivating us to propose the FedMR.
4.2. Server-Side Multimodal Encoding
To efficiently incorporate multimodal item content while mitigating client resource limitations and communication overhead, FedMR performs feature encoding centrally on the server. Leveraging a pre-trained FM $h(\cdot)$, FedMR encodes raw visual features $\mathbf{E}_{v}$ and textual contents $\mathbf{E}_{c}$ into dense semantic embeddings:
(3) $\mathbf{V}=h(\mathbf{E}_{v}),\quad\mathbf{C}=h(\mathbf{E}_{c}).$
This server-side encoding yields communication-efficient, high-quality multimodal representations. Alongside these content embeddings $\mathbf{V}$ and $\mathbf{C}$, FedMR maintains the standard item ID embedding matrix $\mathbf{D}$ to capture collaborative filtering signals (Liu et al., 2023). The three representations $\mathbf{V}$, $\mathbf{C}$ and $\mathbf{D}$ then used to form the input for the subsequent fusion module of FedMR.
4.3. Mixing Feature Fusion Module
FedMR introduce a novel Mixing Feature Fusion Module (MFFM), designed to dynamically fuse the server-provided global representations with client-specific context. MFFM operates in two stages:
Server-Side Multi-Strategy Fusion.
To capture diverse feature interactions, the server employs a predefined set of $|\mathcal{G}|$ distinct fusion strategies $\{g_{j}\}_{j=1}^{|\mathcal{G}|}$. Each strategy $g_{j}$, parameterized by $\gamma_{j}$, combines the multimodal embeddings $\mathbf{V}$ and $\mathbf{C}$, and the ID embedding $\mathbf{D}$ to generate a specific fused representation $\mathbf{F}_{j}$:
(4) $\mathbf{F}_{j}=g_{j}(\mathbf{V},\mathbf{C},\mathbf{D};\gamma_{j}).$
Using multiple strategies produces a diverse set of feature representations $\{\mathbf{F}_{j}\}_{j=1}^{|\mathcal{G}|}$, offering various perspectives on item characteristics. These representations are then transmitted to the clients.
Client-Side Personalized Refinement.
To address user heterogeneity, each client $u$ employs a personalized refinement mechanism. A user-specific router module $\phi_{u}$, parameterized by local parameters $\varphi_{u}$, dynamically computes importance weights $\mathbf{w}_{u}$ for the received server representations based on the client’s interaction history $\mathbf{r}_{u}$:
(5) $\mathbf{w}_{u}=\phi_{u}(\{\mathbf{F}_{j}\}_{j=1}^{|\mathcal{G}|};\mathbf{r}_{u},\varphi_{u}),$
where $\mathbf{w}_{u}\in\mathbb{R}^{|\mathcal{G}|}$ is typically normalized. Each weight $w_{u,j}$ signifies the relevance of the $j$-th feature perspective $\mathbf{F}_{j}$ for user $u$. The final personalized item representation $\bar{\mathbf{F}}^{(u)}$ is obtained by adaptively combining the server representations using these learned weights:
(6) $\bar{\mathbf{F}}^{(u)}=\sum_{j=1}^{|\mathcal{G}|}w_{u,j}\mathbf{F}_{j}.$
This adaptive weighting, learned locally based on $\mathbf{r}_{u}$, allows FedMR to tailor global representations to individual user preferences and local data characteristics, directly mitigating the impact of non-IID data and enhancing personalization with privacy-preserving.
4.4. Objective of FedMR
By substituting the representation $\bar{\mathbf{F}}_{i}^{(u)}$ corresponding to item $i$ for the standard ID embedding $\mathbf{D}_{i}$ within the prediction function $f$ in Eq. (2), the overall objective in Eq. (1) becomes the joint optimization of global and local parameters to obtain the predicted scores $\hat{r}_{u,i}$:
(7) $\min_{\{\theta_{u}\},\{\varphi_{u}\},\{\gamma_{j}\},\mathbf{D}}\sum_{u\in\mathcal{U}}\alpha_{u}\sum_{(u,i)\in\mathcal{O}}\mathcal{L}_{\text{recon}}(f(\bar{\mathbf{F}}_{i}^{(u)};\theta_{u}),r_{u,i}).$
Eq. (7) maintains compatibility with existing FedRec frameworks using prediction functions like $f$, allowing them to benefit from richer, personalized, multimodal item representations while preserving privacy. This leads to a more comprehensive understanding of item features and enhances the model’s grasp of user preferences.
Figure 1. The framework of FedMR. It comprises two components: (1) Server-Side Multi-Strategy Fusion, which fuses multimodal features on server, and (2) Client-Side Personalized Refinement, which dynamically adjusts representations based on user finteractions for personalized recommendations on each client while preserving user’s data privacy.
Algorithm 1 FedMR
Input: $\mathbf{R}$, $\eta$, $T$, $A$
Initialize: $\{\theta_{u}\}$, $\{\varphi_{u}\}$, $\{\gamma_{j}\}$, $\mathbf{D}$
MultiStrategyFusion:
1:$\mathbf{V}\leftarrow h(\mathbf{E}_{v})$, $\mathbf{C}\leftarrow h(\mathbf{E}_{c})$ according to Eq. (3);
2:for $t=1,2,\dots,T$ do
3:    $S_{t}\leftarrow$ randomly select $n_{s}$ from $n$ clients;
4:    Compute $\mathbf{F}_{j}$ according to Eq. (4), $j=1,2,\cdots|\mathcal{G}|$;
5:    for all client index $u\in S_{t}$ do;
6:         $\hat{\mathbf{r}}_{u}$, $\nabla_{\mathbf{D}}^{(u)}$, $\{\nabla^{(u)}_{j}\gamma\}\leftarrow\text{ClientUpdate}(\{\mathbf{F}_{j}\}_{j=1}^{|\mathcal{G}|})$;
7:    end for
8:    $\mathbf{D}\leftarrow\mathbf{D}-\eta\sum_{u\in S_{t}}\alpha_{u}\nabla_{\mathbf{D}}^{(u)}$;
9:    $\gamma_{j}\leftarrow\gamma_{j}-\eta\sum_{u\in S_{t}}\alpha_{u}\nabla^{(u)}_{j}\gamma$, $j=1,2,\cdots,|\mathcal{G}|$;
10:end for
11:return: $\hat{\mathbf{R}}=[\hat{r}_{1},\hat{r}_{2},\dots,\hat{r}_{n}]^{T}$
PersonalizedRefinement:
1:$\nabla_{\mathbf{D}}^{(u)}\leftarrow 0$; $\triangleright$ Initialize gradient accumulators
[…]
3:for $a=1,2,\dots,A$ do
4:    Compute partial gradients $\nabla_{\theta_{u}}$, $\nabla_{\varphi_{u}}$, $\nabla_{\mathbf{D}}$ and $\{\nabla_{\gamma_{j}}\}$ by differentiating Eq. (7);
5:    Update $\theta_{u}\leftarrow\theta_{u}-\eta\nabla_{\theta_{u}}$triangleright Update local prediction parameters;
6:    Update $\varphi_{u}\leftarrow\varphi_{u}-\eta\nabla_{\varphi_{u}}$triangleright Update local router parameters;
7:    Accumulate $\nabla_{\mathbf{D}}^{(u)}\leftarrow\nabla_{\mathbf{D}}^{(u)}+\nabla_{\mathbf{D}}$triangleright Accumulate gradients;
8:    Accumulate $\nabla^{(u)}_{j}{\gamma}\leftarrow\nabla^{(u)}_{j}{\gamma}+\nabla_{\gamma_{j}}$, $j=1,2,\cdots|\mathcal{G}|$;
9:    Compute $\hat{\mathbf{r}}_{u}$ according to Eq. (2);
10:end for
11:return: $\hat{\mathbf{r}}_{u}$, $\nabla_{\mathbf{D}}^{(u)}$, $\{\nabla^{(u)}_{j}{\gamma}\}$
4.5. Unified Framework and Algorithm
The FedMR framework, illustrated in Fig. 1, integrates multimodal information into FedRecs while preserving privacy through an alternating optimization procedure, as shown in Alg. 1). In each communication round $t$, the server first computes the multi-strategy fused representations $\{\mathbf{F}_{j}\}_{j=1}^{|\mathcal{G}|}$ in Eq. (4) using current global parameters ($\mathbf{D},\{\gamma_{j}\}$) and transmits them to a randomly selected client subset $S_{t}$. Each client $u\in S_{t}$ then utilizes its personalized router $\phi_{u}$ in Eq. (5) to compute user-specific weights $\mathbf{w}_{u}$ based on its local history $\mathbf{r}_{u}$, generating the personalized item representation $\bar{\mathbf{F}}^{(u)}$ in Eq. (6). Using $\bar{\mathbf{F}}^{(u)}$, the client performs local updates on its parameters $\theta_{u}$ and $\varphi_{u}$ by minimizing the objective in Eq. (7) on its private data. Concurrently, it computes gradients w.r.t. the global parameters, $\nabla_{\mathbf{D}}^{(u)}$ and $\{\nabla^{(u)}_{\gamma_{j}}\}_{j=1}^{|\mathcal{G}|}$. These gradients are accumulated and sent back to the server. Finally, the server aggregates the received gradients from all clients in $S_{t}$, typically weighted by $\alpha_{u}$, and updates the global parameters $\mathbf{D}$ and $\{\gamma_{j}\}$. This iterative exchange enables collaborative learning. Crucially, the client-side personalized refinement directly addresses data heterogeneity, while privacy is maintained as only gradients are communicated. FedMR’s design thus flexibly integrates multimodal data, enables fine-grained personalization, maintains compatibility with ID-based methods, and ensures efficiency and accuracy within privacy constraints.
4.6. Convergence Analysis
Here, we provide a theoretical analysis of the FedMR under assumptions for non-convex federated optimization. For simplicity, we denote our learnable parameters as $\mathbf{w}=(\Theta_{g},\{\Theta_{l,u}\}_{u\in\mathcal{U}})$, where $\Theta_{g}=\{\mathbf{D},\{\gamma_{j}\}\}$ represents the set of global parameters, and $\Theta_{l,u}=\{\theta_{u},\varphi_{u}\}$ denotes the local parameters for client $u$, capturing user-specific prediction and router parameters. The global objective is:
(8) $F(\mathbf{w})\;=\;\sum_{u\in\mathcal{U}}\alpha_{u}\,F_{u}(\mathbf{w}),$
where each $F_{u}(\mathbf{w})$ is the local objective associated with user $u$, and $\alpha_{u}>0$ with $\sum\alpha_{u}=1$. We aim to find $\mathbf{w}$ that minimizes $F(\mathbf{w})$.
To establish the convergence guarantees, we rely on the following assumptions commonly used in the analysis of FL algorithms:
Assumption 1 (Smoothness).
Each local objective $F_{u}(\mathbf{w})$ is $L$-smooth w.r.t. $\mathbf{w}$. Hence, the global objective $F$ is also $L$-smooth:
(9) $\|\nabla F_{u}(\mathbf{w})-\nabla F_{u}(\mathbf{w}^{\prime})\|\;\leq\;L\,\|\mathbf{w}-\mathbf{w}^{\prime}\|.$
Assumption 2 (Bounded Gradient Variance).
For the stochastic gradient $g_{l,u}(\mathbf{w};\xi)$ of $F_{u}$, we have $\mathbb{E}_{\xi}[g_{l,u}]=\nabla_{\Theta_{l,u}}F_{u}(\mathbf{w})$ and $\mathbb{E}_{\xi}[\|g_{l,u}-\nabla_{\Theta_{l,u}}F_{u}(\mathbf{w})\|^{2}]\leq\sigma_{l}^{2}$. Furthermore, there exists $G>0$ such that $\|\nabla F_{u}(\mathbf{w})\|\leq G$ in expectation across clients.
Assumption 3 (Bounded Heterogeneity).
The variance of all local gradients across clients is bounded as follows:
$\mathbb{E}_{u}\bigl{[}\|\nabla F_{u}(\mathbf{w})-\nabla F(\mathbf{w})\|^{2}\bigr{]}\;\leq\;\sigma^{2}.$
These assumptions ensure that each $F_{u}$ is reasonably well-behaved and client heterogeneity is bounded. Next, we formalize local drift lemmas and provide the main theorem on FedMR’s convergence.
Let $A$ be the local number of SGD steps per communication round. During each round $t$, a subset of clients $\{u\in S_{t}\}$ updates their local parameters $\Theta_{l,u}$ using a local learning rate $\eta$, while the global parameters $\Theta_{g}$ remain fixed until aggregation. We can show:
Lemma 0 (Local Drift Bound).
Under Assumptions 1 and 2, the expected deviation in local parameters after $A$ updates is bounded by a term proportional to $A\eta^{2}(G^{2}+\sigma_{l}^{2})$.
Lemma 1 induces a gradient bias when uploading local updates to the server. Combining the drift bound with smoothness yields:
Lemma 0 (Gradient Bias Bound).
Let $\bar{g}_{g}^{t}$ denote the averaged gradient for global parameters $\Theta_{g}$ after $A$ local updates on the sampled clients, and $B_{t}=\mathbb{E}[\bar{g}_{g}^{t}|\mathbf{w}^{t}]-\nabla_{\Theta_{g}}F(\mathbf{w}^{t})$ be the gradient bias due to local updates relative to the global parameters $\mathbf{w}^{t}$ at the start of the round. Under Assumption 1 and Lemma 1, We then have:
(10) $\mathbb{E}[\|B_{t}\|^{2}]\leq L^{2}C_{1}A\eta^{2}(G^{2}+\sigma_{l}^{2}).$
We now can show FedMR converges to a stationary point of $F(\mathbf{w})$ in a manner consistent with typical non-convex FL rates:
Theorem 3 (Convergence of FedMR).
Suppose Assumptions 13 hold and let $\eta=c_{\eta}/\sqrt{T}$ for a suitable constant $c_{\eta}>0$. Then for $T$ communication rounds, the FedMR algorithm satisfies:
(11) $\small\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\bigl{[}\|\nabla F(\mathbf{w}^{t})\|^{2}\bigr{]}\;\leq\;O\Bigl{(}\frac{1}{\sqrt{T}}\Bigr{)}\;+\;O\Bigl{(}\frac{AG^{2}}{T}\Bigr{)}\;+\;O\Bigl{(}\frac{\sigma^{2}+G^{2}}{\sqrt{T}\,n_{s}}\Bigr{)}.$
Hence, FedMR achieves a convergence rate of $O(T^{-1/2})$ to a stationary point under non-convex settings, matching the standard rates in federated optimization. The above theorem indicates that FedMR’s multimodal feature fusion and personalization do not degrade the asymptotic rate, and it remains comparable to baseline FL methods. The result also quantifies how larger local steps $A$, stronger data heterogeneity $\sigma^{2}$, or fewer sampled clients $n_{s}$ can slow practical convergence. Nonetheless, for moderate $A$, stable $\sigma^{2}$, and typical $n_{s}$, FedMR effectively converges and benefits from the rich multimodal item representations under the federated setup.
4.7. Complexity Analysis
In FedMR, the server initially encodes multimodal features into fixed embeddings $\mathbf{V},\mathbf{C}\in\mathbb{R}^{|\mathcal{I}|\times d}$ using an FM $h(\cdot)$, contributing primarily $O(|\mathcal{I}|d)$ space complexity each, as this is often pre-computed. The server also stores the learnable ID embedding $\mathbf{D}\in\mathbb{R}^{|\mathcal{I}|\times d}$, adding $O(|\mathcal{I}|d)$ space. In each round, the server computes $|\mathcal{G}|$ fused representations $\{\mathbf{F}_{j}\}_{j=1}^{|\mathcal{G}|}$ and aggregates gradients, and the dominant per-round time complexity on the server is $O(|\mathcal{G}||\mathcal{I}|d)$.
On the client side, storing local parameters $K$ and temporarily holding the $|\mathcal{G}|$ received representations requires approximately $O(|\mathcal{I}|d+K)$ space. The client’s per-iteration time complexity for local updates, involving the prediction function $f$ and the MFFM router $\phi_{u}$, is denoted as $O(P)$. Therefore, the total system space complexity is roughly $O(|\mathcal{U}|(|\mathcal{I}|d+K)+|\mathcal{I}|d)$, and the total time complexity per round involves server computation plus aggregated client computation, scaling approximately as $O(|\mathcal{G}||\mathcal{I}|d+n_{s}AP)$ where $n_{s}=|S_{t}|$ and $A$ is the number of local steps. This analysis indicates manageable overhead for incorporating FedMR to achieve personalized and privacy-preserving multimodal recommendation.
4.8. Privacy Preservation
FedMR ensures user privacy primarily through the inherent data localization of the federated learning paradigm, where raw interaction data $\mathbf{r}_{u}$ remains on the client’s device. Communication is limited to model gradients or updates. The MFFM is designed to preserve this standard: server-side computations involve only global representations, while the client-specific router $\phi_{u}$ and its parameters $\varphi_{u}$ remain entirely local. The gradients $\nabla\mathbf{D}^{(u)}$ and $\{\nabla^{(u)}_{j}\gamma\}$ shared with the server are analogous to those shared in standard FedRecs, introducing no additional direct data leakage channels and aligning with current privacy practices (Li et al., 2023a; Zhang et al., 2023c; Li et al., 2024c; Feng et al., 2024; Wu et al., 2024). Privacy is further enhanced by standard techniques like random client sampling in each round (Chai et al., 2020; Li et al., 2023a). Moreover, FedMR is fully compatible with advanced privacy-enhancing technologies such as differential privacy (DP) (Abadi et al., 2016), which can be applied by adding noise to gradients before transmission, and other secure aggregation protocols. Our ablation studies provide empirical evidence that DP can be incorporated without drastically compromising the performance.
5. Experiments
5.1. Datasets
We evaluate FedMR’s performance on four public datasets from the NineRec benchmark [^1] Publicly available at: https://github.com/westlake-repl/NineRec (Zhang et al., 2024a): KU, Bili_Food (Food), Bili_Dance (Dance), and Bili_Movie (Movie). Detailed statistics are provided in Table 1. The KU dataset contains short video interactions from the KuaiShou platform. The Food, Dance, and Movie datasets comprise video interactions from distinct categories on the BiliBili platform. These datasets feature diverse content and user interaction patterns but share the challenge of extreme data sparsity (over 95%). We chose these four datasets from the NineRec benchmark for comprehensive evaluation. They represent diverse real-world scenarios, originating from two distinct platforms and covering varied content domains at different scales. This diversity, coupled with their inherent high sparsity, provides a robust testbed for evaluating FedMR’s adaptability and performance across challenging, realistic conditions. Following standard practice, we filter out users with fewer than 5 interactions to mitigate cold-start issues. For data preprocessing, missing item images are imputed using the average visual features of existing items, and missing item titles are replaced with the placeholder text ”The title is missing.”
Table 1. The statistical information of the datasets used in our work includes the following: #Interactions represents the number of observed interactions. #Users denotes the number of users. #Items indicates the number of items. Sparsity is the percentage of #Interactions out of the total possible ratings.
Dataset #Users #Items #Interactions Sparsity
KU 204 560 3488 96.95%
Dance 10231 1676 80086 99.53%
Food 5990 1125 36482 99.46%
Movie 15908 2528 111091 99.72%
5.2. Baselines
To evaluate the performance of FedMR comprehensively, we compare it with three state-of-the-art centralized MMRecs and five widely used ID-based FedRecs for recommendation tasks with implicit feedback: (1) VBPR (He and McAuley, 2016): Incorporates visual features into personalized ranking to enhance recommendation accuracy for implicit feedback. (2) BM3 (Zhou et al., 2023a): Integrates multimodal features using a self-supervised learning framework. (3) MGCN (Yu et al., 2023): Employs a multi-view graph convolutional network to capture multimodal features of users and items. (4) FedAvg (McMahan et al., 2017): The foundational federated learning method, using weighted aggregation based on local client data sizes. (5) FCF (Ammad-Ud-Din et al., 2019): Combines collaborative filtering with federated learning, performing equal-weighted average aggregation across clients. (6) FedNCF (Perifanis and Efraimidis, 2022): Updates user embeddings locally on clients and synchronizes item embeddings globally on the server. (7) PFedRec (Zhang et al., 2023c): Aggregates client-side item embeddings on the server to enable bilateral personalized recommendations. (8) FedRAP (Li et al., 2023a): Achieves fine-grained personalization for users and items through additive modeling.
The centralized MMRec methods VBPR, BM3 and MGCN serve as benchmarks to gauge the performance potential achievable without federated constraints, highlighting the gap federated approaches aim to close. The selected four federated methods represent common ID-based FedRec approaches. We compare the vanilla versions of the ID-based FedRec baselines against the versions enhanced with FedMR (w/ FedMR) to directly measure the performance gains provided by our framework’s multimodal capabilities [^2] We do not include direct comparisons with AMMFRS (Feng et al., 2024) and FedMMR (Li et al., 2024a) as their official code implementations were not publicly available at the time of our experiments, preventing a fair and reproducible comparison..
5.3. Experimental Setting
Following prior work (Zhang et al., 2023c; Li et al., 2023a), we employ negative sampling during training and use the leave-one-out strategy for validation and testing. Hyperparameters for all baselines are tuned via grid search on the validation set, including the learning rate $\eta$ within $\{10^{i}\mid i=-4,\ldots,-1\}$ and method-specific parameters. We use CLIP (Radford et al., 2021a) as the common foundation model for extracting visual and textual features, implemented via OpenCLIP (Ilharco et al., 2021) with pre-trained ViT-B-32 weights from OpenAI (Radford et al., 2021b; Schuhmann et al., 2022). A linear mapping layer projects multimodal and ID embeddings to a shared latent dimension of 64 for efficiency. The training batch size is 2048. All models incorporating hidden layers use a three-layer MLP structure. For federated methods, clients perform $5$ local training epochs per communication round using a consistent aggregation strategy for both vanilla and FedMR-integrated versions.
5.4. Evaluation
In our experiments, we treat the observed interactions as positive samples and the rest as negative samples. We sort the predicted scores between users and all candidate items in descending order, masking the observed positive samples from the training set. Then, we evaluate the recommendation performance using two widely-used metrics for implicit feedbacks: Hit Rate (HR@K) and Normalized Discounted Cumulative Gain (NDCG@K), which are formally defined in (He et al., 2020). In the experiments, we set $K=50$.
Table 2. Performance comparison of Centralized (Cent) and Federated (Fed) Methods on four real-world datasets using HR@50 and NDCG@50. The federated methods are evaluated in both their original versions (vanilla) and enhanced with FedMR (w/ FedMR). The percentages in parentheses indicate the improvement of the version with FedMR compared to the original version.
KU Dance Food Movie
Method HR@50 NDCG@50 HR@50 NDCG@50 HR@50 NDCG@50 HR@50 NDCG@50
VBPR 0.4412 0.1547 0.2306 0.0747 0.2472 0.0790 0.1429 0.0486
BM3 0.4167 0.1595 0.2654 0.0856 0.2563 0.0820 0.1917 0.0666
Cent MGCN 0.3382 0.1508 0.2363 0.0767 0.2409 0.0777 0.1697 0.0574
vanilla 0.1127 0.0562 0.0782 0.0201 0.0696 0.0191 0.0552 0.0158
FedAvg w/ FedMR 0.5147 0.1491 0.1007 0.0261 0.0948 0.0268 0.0564 0.0160
Impr. 356.70% $\uparrow$ 165.30% $\uparrow$ 28.77% $\uparrow$ 29.85% $\uparrow$ 36.21% $\uparrow$ 40.31% $\uparrow$ 2.17% $\uparrow$ 1.27% $\uparrow$
vanilla 0.2255 0.0643 0.0676 0.0177 0.0820 0.0219 0.0581 0.0163
FCF w/ FedMR 0.5539 0.1597 0.0699 0.0182 0.0918 0.0247 0.0667 0.0180
Impr. 145.63% $\uparrow$ 148.37% $\uparrow$ 3.40% $\uparrow$ 2.75% $\uparrow$ 11.95% $\uparrow$ 12.79% $\uparrow$ 14.80% $\uparrow$ 10.44% $\uparrow$
vanilla 0.1176 0.0249 0.0221 0.0055 0.0548 0.0163 0.0185 0.0054
FedNCF w/ FedMR 0.1961 0.0609 0.0708 0.0187 0.0843 0.0243 0.0406 0.0125
Impr. 66.75% $\uparrow$ 144.58% $\uparrow$ 220.36% $\uparrow$ 240.00% $\uparrow$ 53.83% $\uparrow$ 49.08% $\uparrow$ 119.46% $\uparrow$ 131.48% $\uparrow$
vanilla 0.0637 0.0230 0.0252 0.0068 0.0327 0.0093 0.0170 0.0042
PFedRec w/ FedMR 0.2157 0.0616 0.0392 0.0092 0.0836 0.0207 0.0413 0.0132
Impr. 238.62% $\uparrow$ 167.83% $\uparrow$ 55.56% $\uparrow$ 35.29% $\uparrow$ 155.66% $\uparrow$ 122.58% $\uparrow$ 142.94% $\uparrow$ 214.29% $\uparrow$
vanilla 0.4706 0.1086 0.0682 0.0186 0.0696 0.0191 0.0453 0.0136
Fed FedRAP w/ FedMR 0.4824 0.1305 0.0895 0.0247 0.1142 0.0294 0.0586 0.0199
Impr. 2.51% $\uparrow$ 20.17% $\uparrow$ 31.19% $\uparrow$ 32.91% $\uparrow$ 64.08% $\uparrow$ 53.93% $\uparrow$ 29.36% $\uparrow$ 46.32% $\uparrow$
5.5. Comparison Analysis
The performance comparison detailed in Table 2 reveals key insights into FedMR’s effectiveness relative to both ID-based federated baselines and centralized multimodal systems. The results consistently demonstrate FedMR’s significant value in bridging the gap between privacy-preserving FL and effective multimodal recommendation.
Enhancing ID-based FedRec. FedMR consistently and substantially elevates the performance of all evaluated ID-based FedRecs across the four datasets. The degree of improvement varies depending on the baseline architecture. Performance gains are particularly dramatic for methods initially exhibiting lower performance, such as FedAvg, FedNCF, and PFedRec, especially on datasets where multimodal signals prove highly beneficial like KU and Food. This trend suggests FedMR effectively integrates crucial multimodal context and personalized fusion capabilities that compensate for the limitations of simpler ID-based approaches in content understanding and heterogeneity handling. While FedMR also enhances more sophisticated baselines like FedRAP, the relative improvement tends to be more moderate, indicating diminishing returns when the base model already incorporates strong personalization mechanisms, though the added multimodal dimension still provides value.
Performance Relative to Centralized Methods. When compared against centralized MMRec benchmarks, FedMR-enhanced methods demonstrate encouraging results, significantly narrowing the typical performance gap observed between federated and centralized systems. Notably, on the KU dataset, FedMR integrated with FedAvg and FCF achieves performance surpassing the centralized counterparts. This suggests that FedMR’s adaptive client-side fusion can be highly effective, potentially adapting better to user nuances than global centralized fusion in certain scenarios. However, a performance gap often persists on the larger Dance, Food, and Movie datasets compared to the best centralized methods. This trend indicates that while FedMR offers substantial improvements, the inherent challenges under the federated recommmendation settings, such as data fragmentation and privacy constraints limiting information sharing, still make fully matching centralized performance difficult, particularly on larger-scale, complex datasets. Furthermore, the lower absolute scores achieved by all methods on datasets like Dance and Movie suggest these represent intrinsically more challenging recommendation tasks, possibly due to complex user behavior or less discriminative multimodal features.
Summary. In summary, FedMR’s integration of multimodal information through the server-client architecture and the adaptive MFFM consistently improves upon diverse ID-based FedRecs. FedMR significantly narrows, and sometimes closes or even surpasses, the performance gap relative to centralized MMRecs. The remaining gap on more challenging datasets highlights the ongoing need for advancements in FedMMRec. FedMR represents a crucial step in this direction, showcasing the effectiveness of adaptive, personalized fusion mechanisms in unlocking multimodal data’s potential under privacy-preserving federated constraints. Specifically, its ability to dynamically adapt to user preferences underscores its value for developing robust and personalized FedMMRecs.
5.6. Ablation Study
To comprehensively evaluate FedMR’s contributions and key design choices, we conduct ablation studies analyzing four crucial aspects. We investigate the effectiveness of the MFFM compared to simpler alternatives, assess the individual contributions of visual, textual, and ID modalities, examine the impact of client heterogeneity related to user activity levels, and evaluate the model’s robustness when incorporating noise-based privacy enhancement. These analyses collectively highlight the significance of FedMR’s adaptive multimodal fusion and illustrate the interplay between modality utilization, user diversity, and privacy considerations in FedRec.
Figure 2. Performance comparison of different feature fusion strategies in various FedRecs combined with FedMR on KU.
5.6.1. Ablation Study on Fusion Strategy
Fig. 2 shows the HR@50 and NDCG@50 results for five FedRec methods with different fusion strategies on the KU dataset. The results highlight the critical role of fusion strategies in performance. Among them, the Mix strategy, which dynamically combines fusion methods based on user interaction history, achieves the best performance. In contrast, the Sum strategy performs the worst across all metrics. The Sum strategy, which simply adds features, fails to account for complex interactions among modalities, often leading to information loss and poor handling of long-tail distributions. In contrast, the MLP strategy better captures nonlinear interactions but is limited by its network depth and parameter count. The Gate strategy dynamically learns feature weights, optimizing the importance of different modalities. However, the Mix strategy combines all fusion methods, dynamically adjusting their contributions, enabling it to generate fine-grained personalized item representations and better capture complex relationships among features. This study underscores the importance of personalized item representations in FedMMRec. By leveraging the MFFM, FedMR allows each client to dynamically adjust fusion weights based on user interaction history, enabling fine-grained personalization and enhancing recommendation accuracy while maintaining privacy under the federated settings.
Table 3. Performance comparison of FedRecs with and without multimodal and ID features on KU.
Method Metric FedMR w/o V w/o C w/o D
FedAvg HR@50 0.5147 0.4069 (20.94% $\downarrow$) 0.3725 (27.63% $\downarrow$) 0.5094 (1.03% $\downarrow$)
NDCG@50 0.1491 0.1391 (6.71% $\downarrow$) 0.1287 (13.68% $\downarrow$) 0.1407 (5.66% $\downarrow$)
FCF HR@50 0.5539 0.2843 (48.67% $\downarrow$) 0.4216 (23.89% $\downarrow$) 0.2353 (57.52% $\downarrow$)
NDCG@50 0.1597 0.0748 (53.16% $\downarrow$) 0.0965 (39.57% $\downarrow$) 0.0578 (63.81% $\downarrow$)
FedNCF HR@50 0.1961 0.1569 (19.99% $\downarrow$) 0.1863 (5.00% $\downarrow$) 0.1602 (18.31% $\downarrow$)
NDCG@50 0.0609 0.0431 (29.23% $\downarrow$) 0.0541 (11.17% $\downarrow$) 0.0538 (11.66% $\downarrow$)
PFedRec HR@50 0.2157 0.1990 (7.74% $\downarrow$) 0.1941 (10.01% $\downarrow$) 0.1794 (16.83% $\downarrow$)
NDCG@50 0.0616 0.0469 (23.86% $\downarrow$) 0.0537 (12.82% $\downarrow$) 0.0473 (23.21% $\downarrow$)
FedRAP HR@50 0.3824 0.2647 (30.78% $\downarrow$) 0.3235 (15.40% $\downarrow$) 0.1863 (51.29% $\downarrow$)
NDCG@50 0.1305 0.0762 (41.61% $\downarrow$) 0.0714 (45.29% $\downarrow$) 0.0490 (62.43% $\downarrow$)
5.6.2. Ablation Study on Modality Efficiency
We assess the contribution of visual, textual, and ID feature types by evaluating FedMR-enhanced model performance upon removing each source individually, referencing the updated results in Table 3. The findings reveal a clear interdependence among these modalities, confirming their complementary roles. ID features remain foundational for personalization, and their removal consistently causes the most significant performance degradation across baseline methods, often proving the most critical factor, particularly impacting models with strong personalization components like FedRAP. Visual features are also highly influential, with their absence substantially hindering performance, especially for visually-reliant architectures such as FCF. Textual features provide valuable semantic context, and removing them leads to notable accuracy decreases for various models, confirming their complementary contribution beyond visual and ID signals. Crucially, we can see the degree of reliance on each modality differs markedly across the underlying FedRec architectures. This differential dependency and demonstrated complementarity across modalities necessitate adaptive fusion mechanisms. FedMR’s MFFM addresses this requirement, ensuring robust performance by effectively utilizing all available information sources.
(a) User Distribution
(b) Modality Preferences
(c) Heterogeneity Distribution
(d) Performance Trend
Figure 3. Analysis of user characteristics, heterogeneity, and performance for FedRAP w/ FedMR on the KU dataset. (a) Distribution of users across interaction counts; (b) Variation in user preferences across modalities relative to the global average; (c) Distribution of user heterogeneity, which deviation from global preference, across interaction groups; (d) Recommendation performance trend on Recall@50 and NDCG@50 by user interaction group.
5.6.3. Ablation Study on Client Heterogeneity
We analyze client heterogeneity and its performance implications using the FedRAP enhanced by FedMR on the KU dataset, with results consolidated in Fig. 3. Fig. 3(a) shows that the dataset comprises users with varying activity levels. A key finding is the significant inter-user heterogeneity in modality preferences, as shown in Fig. 3(b). While preferences related to ID features exhibit relatively low variance, user preferences for text and especially vision features vary considerably more, highlighting the diverse ways users perceive and value different types of multimodal information, posing a challenge for standard fusion approaches. Furthermore, as we can see in Fig. 3(c), analyzing the overall heterogeneity distribution reveals that its variance tends to increase with user activity. More active users, therefore, not only deviate from the average preference but also exhibit greater diversity amongst themselves compared to less active users, suggesting that experienced users often develop more distinct and varied multimodal preference patterns, consistent with prior work (Liu et al., 2019; Guo et al., 2024).
This observed trend of increasing heterogeneity variance among more active users directly correlates with the recommendation performance shown in Fig. 3(d). Recall@50 and NDCG@50 tend to decrease for user groups with higher interaction counts. The heightened preference diversity within these active groups makes accurate Top-K prediction intrinsically more difficult. Capturing the varied, potentially niche multimodal preferences of highly heterogeneous users presents a significant challenge, contributing to the observed degradation in average performance metrics for these groups. This challenge underscores the limitations of non-adaptive fusion methods and highlights the core contribution of FedMR. Our framework introduces the MFFM, an innovative client-side mechanism specifically designed for dynamic, personalized adaptation of multimodal representations. By allowing each client to tailor the fusion process based on individual interaction history, MFFM directly addresses the crucial challenge of modality preference heterogeneity observed empirically, offering a vital step towards effective personalization in diverse federated multimodal environments.
Table 4. Performance comparison of FedRec methods combined with FedMR, with and without Gaussian noise (mean=0, variance=0.1) on KU dataset for privacy protection.
Method Metric FedMR w/ noise Degrade
FedAvg HR@50 0.5147 0.4969 3.46% $\downarrow$
NDCG@50 0.1491 0.1416 5.03% $\downarrow$
FCF HR@50 0.5539 0.5378 2.91% $\downarrow$
NDCG@50 0.1597 0.1402 12.21% $\downarrow$
FedNCF HR@50 0.1961 0.1680 14.33% $\downarrow$
NDCG@50 0.0609 0.0420 31.03% $\downarrow$
PFedRec HR@50 0.2157 0.1965 8.90% $\downarrow$
NDCG@50 0.0616 0.0586 4.87% $\downarrow$
FedRAP HR@50 0.3824 0.3725 2.59% $\downarrow$
NDCG@50 0.1305 0.1236 5.29% $\downarrow$
5.6.4. Ablation Study on Privacy Preservation
A key consideration in FedRec is compatibility with privacy-enhancing technologies. We assess FedMR’s robustness by evaluating the performance impact of adding Gaussian noise to client gradients before server aggregation, a common technique related to differential privacy. Table 4 presents these results on the KU dataset. As anticipated, introducing noise creates a privacy-utility trade-off, generally leading to decreased recommendation accuracy across all tested FedMR-enhanced models. However, the extent of this performance degradation differs significantly depending on the base FedRec architecture. Models with simpler structures or strong adaptive components, specifically FedAvg, FCF, and FedRAP, demonstrate considerable resilience, experiencing relatively minor performance reductions. Conversely, methods employing more complex neural networks, such as FedNCF, exhibit higher sensitivity to the injected noise, resulting in a more substantial drop in accuracy. This differential impact suggests that noise may interfere more significantly with the intricate parameter relationships learned by complex models. Nevertheless, the performance reduction remains moderate for most configurations, indicating that FedMR effectively integrates with various baseline methods and maintains reasonable utility under standard noise-based privacy mechanisms, confirming its practical applicability for privacy-conscious recommendation tasks in federated environments.
6. Conclusion
Motivated by privacy concerns and regulations that necessitate alternatives to data-intensive centralized multimodal recommendation, this paper introduced FedMR, a novel framework addressing the critical challenge of effectively integrating multimodal data within privacy-preserving Federated Recommendation Systems. FedMR’s core innovation lies in the client-side Mixing Feature Fusion Module, which facilitates dynamic, personalized adaptation by fusing server-provided multimodal representations with ID-based signals according to individual user interaction histories, explicitly modeling and addressing user heterogeneity in multimodal preference fusion directly on the client. This adaptive mechanism effectively handles diverse user preferences across modalities. Extensive experiments demonstrate that FedMR seamlessly enhances various ID-based FedRecs, substantially boosting their performance and significantly narrowing the gap towards centralized multimodal systems, occasionally even surpassing them. By enabling adaptive multimodal fusion under federated constraints, FedMR offers a practical and robust solution, advancing the state-of-the-art towards effective, personalized, and privacy-preserving multimodal recommendation in federated environments.
Appendix A Algorithm Optimization
A.1. Derivation of the Update for $\gamma_{j}^{(u)}$
The update process for $\{\gamma_{j}^{(u)}\}_{j=1}^{|\mathcal{G}|}$ in Alg. 1 is based on gradient descent. The local update at each client $u$ during the iteration $t$ follows the following rule:
(12) $\gamma_{j}^{(t)}=\gamma_{j}^{(t-1)}-\eta\nabla_{j}^{(u)}\gamma^{(t)},$
where $\eta$ is the learning rate, and $\nabla_{j}^{(u)}\gamma^{(t)}$ represents the gradient of $\gamma_{j}^{(u)}$ at the iteration $t$ on client $u$.
To capture how $\gamma_{j}^{(u)}$ evolves over time, we can unroll this equation over several iterations. By recursively applying the update rule in Eq. (12), we obtain the following:
(13) $\gamma_{j}^{(t)}=\gamma_{j}^{(0)}-\eta\sum_{i=1}^{t}\nabla_{j}^{(u)}\gamma^{(i)},$
where $\gamma_{j}^{(0)}$ is the initial value of $\gamma_{j}^{(u)}$, and the second term is the sum of all gradients computed up to the $t$-th iteration. This expression shows that the updated value of $\gamma_{j}^{(u)}$ depends on its initial value and the accumulated gradient over $t$ iterations.
A.2. Accumulated Gradient
The accumulated gradient $\nabla_{j}^{(u)}\gamma$ at the client $u$ is the sum of the individual gradients across the iterations, which be expressed as:
(14) $\nabla_{j}^{(u)}\gamma=\frac{1}{\eta}\left(\gamma_{j}^{(0)}-\gamma_{j}^{(t)}\right).$
Eq. (14) represents the total change in $\gamma_{j}^{(u)}$ during the local training process. It relates the initial value $\gamma_{j}^{(0)}$ to the current value $\gamma_{j}^{(t)}$, scaled by the learning rate $\eta$.
A.3. Global Aggregation
In federated learning, the global model is updated by aggregating the local updates from multiple clients. For the parameter $\gamma_{j}$, this is achieved by summing the accumulated gradients $\nabla_{j}^{(u)}\gamma$ from all selected clients. The global update for $\gamma_{j}$ can be written as:
(15) $\gamma_{j}=\gamma_{j}-\eta\sum_{u\in\mathcal{S}_{t}}\alpha_{u}\nabla_{j}^{(u)}\gamma,$
where $\mathcal{S}_{t}$ is the set of clients selected during round $t$. This step updates the global parameter $\gamma_{j}$ by aggregating the contributions from each client’s local updates. The derived equations demonstrate that the update process is consistent with the principles of gradient-based optimization. This ensures that $\gamma_{j}^{(u)}$ is updated in an efficient and mathematically sound manner within the FL framework.
Appendix B Theoretical Analysis
B.1. Convergence Analysis
Proof of Lemma 1 (Local Update Bound).
Using the local update rule:
(16) $\Theta_{l,u}^{t,k}=\Theta_{l,u}^{t,k-1}-\eta\,g_{l,u}(\Theta_{l,u}^{t,k-1};\xi),$
and Assumption 2, we have:
(17) $\begin{split}\mathbb{E}\Bigl{\|}\Theta_{l,u}^{t,k}-\Theta_{l,u}^{t}\Bigr{\|}^{2}&=\mathbb{E}\left\|\sum_{i=1}^{k}\eta\,g_{l,u}(\Theta_{l,u}^{t,i-1};\xi)\right\|^{2}\\[4.0pt] &\leq k\,\eta^{2}\sum_{i=1}^{k}\mathbb{E}\left\|g_{l,u}(\Theta_{l,u}^{t,i-1};\xi)\right\|^{2}\quad\text{(by Jensen's inequality)}\\[4.0pt] &\leq k^{2}\,\eta^{2}(G^{2}+\sigma_{l}^{2})\quad\text{(by Assumption 2).}\end{split}$
Jensen’s inequality is used because the square of a sum under an expectation can be bounded by $k$ times the sum of squares.
For $\eta=\tfrac{1}{L\sqrt{A}}$ and $k=A$, it follows that:
(18) $\begin{split}\mathbb{E}\Bigl{\|}\Theta_{l,u}^{t,A}-\Theta_{l,u}^{t}\Bigr{\|}^{2}&\leq A^{2}\,\eta^{2}(G^{2}+\sigma_{l}^{2})=\frac{A(G^{2}+\sigma_{l}^{2})}{L^{2}}.\end{split}$
Defining $C_{1}=A$, we can equivalently express the bound as the followings:
(19) $\begin{split}\mathbb{E}\Bigl{\|}\Theta_{l,u}^{t,A}-\Theta_{l,u}^{t}\Bigr{\|}^{2}\leq C_{1}\,A\,\eta^{2}(G^{2}+\sigma_{l}^{2}).\end{split}$
Proof of Lemma 2 (Gradient Bias Bound).
The gradient bias is given by
(20) $\begin{split}\|B_{t}\|&=\Bigl{\|}\mathbb{E}_{u}\Bigl{[}\nabla_{\Theta_{g}}F_{u}\bigl{(}\Theta_{g}^{t},\Theta_{l,u}^{t,A}\bigr{)}\Bigr{]}-\nabla_{\Theta_{g}}F(\Theta_{g}^{t})\Bigr{\|}\\[4.0pt] &=\Bigl{\|}\mathbb{E}_{u}\Bigl{[}\nabla_{\Theta_{g}}F_{u}\bigl{(}\Theta_{g}^{t},\Theta_{l,u}^{t,A}\bigr{)}-\nabla_{\Theta_{g}}F_{u}\bigl{(}\Theta_{g}^{t},\Theta_{l,u}^{t}\bigr{)}\Bigr{]}\Bigr{\|}\\[4.0pt] &\leq\mathbb{E}_{u}\Bigl{\|}\nabla_{\Theta_{g}}F_{u}\bigl{(}\Theta_{g}^{t},\Theta_{l,u}^{t,A}\bigr{)}-\nabla_{\Theta_{g}}F_{u}\bigl{(}\Theta_{g}^{t},\Theta_{l,u}^{t}\bigr{)}\Bigr{\|}\\[4.0pt] &\leq L\,\mathbb{E}_{u}\Bigl{\|}\Theta_{l,u}^{t,A}-\Theta_{l,u}^{t}\Bigr{\|},\end{split}$
where the last inequality follows from Assumption 1. Squaring both sides and applying the result from Lemma 1 yields:
(21) $\begin{split}\|B_{t}\|^{2}\leq L^{2}\,\mathbb{E}_{u}\Bigl{\|}\Theta_{l,u}^{t,A}-\Theta_{l,u}^{t}\Bigr{\|}^{2}\leq L^{2}\,C_{1}\,A\,\eta^{2}(G^{2}+\sigma_{l}^{2}).\end{split}$
Lemma 0 (Aggregated Gradient Second Moment Bound).
Under Assumptions 2, 3, and Lemma 2, assuming uniform client sampling on $n_{s}$ clients, we can derive the aggregated gradient second moment bound as followings:
(22) $\mathbb{E}[\|\bar{g}_{g}^{t}\|^{2}]\leq 2\mathbb{E}[\|\nabla F(\Theta_{g}^{t})\|^{2}]+2\mathbb{E}[\|B_{t}\|^{2}]+\frac{2(\sigma^{2}+G^{2})}{n_{s}}$
Proof of Lemma 1 (Aggregated Gradient Second Moment).
When Assumptions 2 and 3 hold, we can use the variance decomposition to derive the followings:
(23) $\begin{split}\mathbb{E}\Bigl{\|}\bar{g}_{g}^{t}\Bigr{\|}^{2}&=\Bigl{\|}\mathbb{E}\bar{g}_{g}^{t}\Bigr{\|}^{2}+\mathbb{E}\Bigl{\|}\bar{g}_{g}^{t}-\mathbb{E}\bar{g}_{g}^{t}\Bigr{\|}^{2}\\[3.0pt] &=\Bigl{\|}\nabla F(\Theta_{g}^{t})+B_{t}\Bigr{\|}^{2}+\operatorname{Var}\left(\bar{g}_{g}^{t}\right)\\[3.0pt] &\leq 2\Bigl{\|}\nabla F(\Theta_{g}^{t})\Bigr{\|}^{2}+2\|B_{t}\|^{2}+\frac{1}{n_{s}}\,\mathbb{E}_{u}\Bigl{\|}\nabla_{\Theta_{g}}F_{u}\bigl{(}\Theta_{g}^{t},\Theta_{l,u}^{t,A}\bigr{)}\Bigr{\|}^{2}\\[3.0pt] &\leq 2\Bigl{\|}\nabla F(\Theta_{g}^{t})\Bigr{\|}^{2}+2\|B_{t}\|^{2}+\frac{G^{2}+\sigma^{2}}{n_{s}}.\end{split}$
The last step is based on Assumptions 2 and 3. ∎
Proof of Theorem 1 (Convergence Rate).
Starting from the $L$-smoothness of $F$, which is in Assumpution 1, we have:
(24) $\begin{split}F(\Theta_{g}^{t+1})&\leq F(\Theta_{g}^{t})+\bigl{\langle}\nabla F(\Theta_{g}^{t}),\Theta_{g}^{t+1}-\Theta_{g}^{t}\bigr{\rangle}+\frac{L}{2}\Bigl{\|}\Theta_{g}^{t+1}-\Theta_{g}^{t}\Bigr{\|}^{2}\\[3.0pt] &=F(\Theta_{g}^{t})-\eta\,\bigl{\langle}\nabla F(\Theta_{g}^{t}),\bar{g}_{g}^{t}\bigr{\rangle}+\frac{L\eta^{2}}{2}\Bigl{\|}\bar{g}_{g}^{t}\Bigr{\|}^{2}.\end{split}$
Taking expectations and using Lemmas 2 and 3, we obtain:
(25) $\begin{split}\mathbb{E}F(\Theta_{g}^{t+1})&\leq\mathbb{E}F(\Theta_{g}^{t})-\eta\,\mathbb{E}\Bigl{\langle}\nabla F(\Theta_{g}^{t}),\nabla F(\Theta_{g}^{t})+B_{t}\Bigr{\rangle}+\frac{L\eta^{2}}{2}\,\mathbb{E}\Bigl{\|}\bar{g}_{g}^{t}\Bigr{\|}^{2}\\[3.0pt] &\leq\mathbb{E}F(\Theta_{g}^{t})-\frac{\eta}{2}\,\mathbb{E}\Bigl{\|}\nabla F(\Theta_{g}^{t})\Bigr{\|}^{2}+\frac{\eta}{2}\,\mathbb{E}\|B_{t}\|^{2}\\[3.0pt] &\quad+L\eta^{2}\,\mathbb{E}\Bigl{\|}\nabla F(\Theta_{g}^{t})\Bigr{\|}^{2}+L\eta^{2}\,\mathbb{E}\|B_{t}\|^{2}+\frac{L\eta^{2}(G^{2}+\sigma^{2})}{2n_{s}}.\end{split}$
Here we see the interplay of the gradient norm $\|\nabla F(\Theta_{g}^{t})\|$ and the bias term $B_{t}$. We control these via the learning rate $\eta$. Choosing $\eta\leq\frac{1}{4L}$ ensures certain terms are upper-bounded by $\frac{1}{2}\|\nabla F(\Theta_{g}^{t})\|^{2}$.
For $\eta\leq\frac{1}{4L}$, collecting like terms yields:
(26) $\begin{split}\frac{\eta}{4}\,\mathbb{E}\Bigl{\|}\nabla F(\Theta_{g}^{t})\Bigr{\|}^{2}&\leq\mathbb{E}\Bigl{[}F(\Theta_{g}^{t})-F(\Theta_{g}^{t+1})\Bigr{]}+\frac{5L^{3}C_{1}A\,\eta^{3}(G^{2}+\sigma_{l}^{2})}{2}\\ &\quad+\frac{L\eta^{2}(G^{2}+\sigma^{2})}{2n_{s}}.\end{split}$
Summing over $t=0$ to $T-1$ and setting $\eta=\frac{1}{L\sqrt{T}}$, we obtain
(27) $\begin{split}\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\Bigl{\|}\nabla F(\Theta_{g}^{t})\Bigr{\|}^{2}&\leq\frac{4L\,(F_{0}-F^{*})}{\sqrt{T}}+\frac{10C_{1}AG^{2}}{T}+\frac{2(G^{2}+\sigma^{2})}{\sqrt{T}\,n_{s}}.\end{split}$
The additional term $\mathcal{O}\bigl{(}G^{2}/T\bigr{)}$ is dominated by $\mathcal{O}\bigl{(}G^{2}/\sqrt{T}\bigr{)}$, which concludes the proof. ∎
Table 5. Changes of parameter numbers in FedRec with and without FedMR on the KU Dataset, Compared to Centralized Recommendation Systems.
Method KU Dance Food Movie
Cent VBPR 127552 1482432 904320 2263616
BM3 1752576 8600576 5582848 12815872
MGCN 708864 2564800 1693888 3855104
Fed FedAvg Vanilla 35905 107329 72065 161857
FedMR 455620 1027012 744900 1463236
FCF Vanilla 35905 107329 72065 161857
FedMR 455620 1027012 744900 1463236
FedNCF Vanilla 108737 1535041 921665 1182689
FedMR 564292 2561988 1666500 3833924
PFedRec Vanilla 35905 107329 72065 161857
FedMR 455620 1027012 744900 1463236
FedRAP Vanilla 71745 214593 144065 323649
FedMR 742340 1885124 1320900 2757572
Appendix C Parameter Analysis
Table 5 illustrates the change in the number of learnable parameters on a single client in the KU dataset when FedMR is integrated into FedRec or not, compared to centralized recommendation systems. As shown in the table, the introduction of FedMR significantly increases the parameter size. This is because FedMR incorporates multimodal information (e.g., visual and textual features) alongside ID embeddings, increasing the model’s complexity and consequently requiring more parameters to learn.
Despite the increase in the number of learnable parameters, Table 2 of the main paper demonstrates that FedMR significantly improves the recommendation performance of FedRec on the KU dataset. By combining multimodal features with ID embeddings, FedMR better leverages users’ interaction histories to capture the fine-grained personality of users. Additionally, the mixing feature fusion module in FedMR adjusts the weights of multimodal features based on users’ interaction history, greatly enhancing both the accuracy and personalization of recommendations. Moreover, compared to traditional FedRec methods, which rely solely on the ID embeddings, FedMR enables personalized modeling by integrating multimodal information, thereby further improving recommendation quality. While the model’s complexity increases, this trade-off results in significant performance gains.
As privacy protection is crucial in the federated settings, FedMR processes multimodal data locally on the client and only uploads essential parameters, e.g., ID embeddings and gradients, to the server, avoiding the exposure of raw user data. Although this adds computational overhead and increases parameter size on the client, these sacrifices are necessary to ensure privacy while also boosting recommendation performance. Thus, the increased learnable parameter size is not only a necessary cost for improved recommendation performance but also a vital measure for ensuring privacy.
Appendix D More Experiments
D.1. More Experiment Details
All baseline methods are implemented in PyTorch (Paszke et al., 2019), and their official implementations are publicly available. Experiments are conducted on a machine with a 3.60 GHz Intel Xeon Gold 6244 CPU, three Quadro RTX 6000 GPUs, and 395 GB of RAM.
Figure 4. Training loss progression over epochs for baseline methods enhanced with FedMR on the KU dataset, illustrating convergence behavior.
D.1.1. Convergence Analysis
To assess the training stability and convergence properties of the baseline methods when enhanced with FedMR, we plot their training loss progression over 50 epochs on the KU dataset in Figure 4. The loss curves demonstrate that all FedMR-enhanced models exhibit stable convergence behavior. The training loss decreases consistently across epochs for all methods, with a sharp reduction typically observed in the initial epochs followed by more gradual convergence towards a stable value. This indicates that the FedMR framework integrates well with diverse base FedRec architectures, allowing for effective optimization and stable learning dynamics. While the final converged loss values vary slightly across methods, all models successfully minimize the training objective, confirming the robustness of the training process when incorporating FedMR’s multimodal fusion capabilities.
Figure 5. Sensitivity analysis of 4 FedMR-enhanced baselines (FCF, FedAvg, FedNCF, PFedRec) on the KU dataset. Performance is plotted against varying learning rates (lr)and L2 regularization (l2_reg) strengths.
Figure 6. Hyperparameter tuning for FedMR-enhanced FedRAP on the KU dataset. Heatmaps visualize HR@50 and NDCG@50 across personalization parameters $\alpha$ and $\beta$ for different learning rates (lr) and L2 regularization (l2_reg). Performance intensity is indicated by color.
D.1.2. Hyperparameter Analysis
We analyze the sensitivity of FedMR-enhanced models to key hyperparameters, specifically the learning rate and L2 regularization strength, using the KU dataset. Fig. 5 illustrates this for FedMR applied to FCF, FedAvg, FedNCF, and PFedRec. Across these models, the learning rate exhibits a significant impact, with performance generally peaking at intermediate values and decreasing notably at lower or higher rates. This highlights the necessity of selecting an appropriate learning rate for effective convergence. The influence of L2 regularization appears moderate in comparison, though still relevant; no single $l2\_reg$ value is universally optimal, with the best performance achieved at different strengths depending on the base model and learning rate, necessitating model-specific tuning.
The tuning process for the FedMR-enhanced FedRAP model, which includes its specific personalization parameters $\alpha$ and $\beta$, is detailed in Fig. 6. The heatmaps clearly indicate high sensitivity to the choices of $\alpha$ and $\beta$, with distinct optimal regions emerging. Furthermore, these optimal regions shift based on the learning rate, signifying a strong interaction between $lr$, $\alpha$, and $\beta$. This underscores the need for careful co-tuning, typically via grid search, to identify the best configuration for FedMR-enhanced FedRAP. Overall, while the models demonstrate reasonable robustness, achieving optimal results requires careful hyperparameter tuning, particularly for the learning rate and any model-specific parameters. The best hyperparameters identified through validation during this analysis informed the settings used for the main experimental results.
Figure 7. Visualization of multimodal features and ID features mapped into a shared representation space.
D.1.3. Ablation Study on Modality Distributions
To analyze the interaction between multimodal features and ID features, we use FedRAP as the backbone for FedMR. Fig. 7 visualizes the distribution of these features on the KU dataset. The first subfigure shows that globally shared ID features and multimodal features occupy distinct subregions in the shared representation space, a phenomenon known as the ”Modality Gap” (Liang et al., 2022; Zhang et al., 2023a). This separation enhances the robustness of multimodal data processing by reducing noise interference. The subsequent subfigures illustrate how personalized ID features influence this distribution. Incorporating personalized ID embeddings (e.g., for Users 00 and $42$) causes the ID features to expand toward the regions of visual and textual features, effectively shrinking the modality gap. This adjustment facilitates multimodal fusion and improves the accuracy of personalized recommendations. By comparing users, it is evident that the model dynamically adapts feature distributions based on user-specific behaviors and preferences, better addressing individual needs. This demonstrates the critical role of personalized ID features in federated recommendation systems, enabling effective personalization while preserving user privacy.
Figure 8. Performance comparsions of different FedRec methods with FedMR under varying client sampling ratios on HR@50 and NDCG@50.
D.2. Ablation Study on Client Sampling
To investigate the impact of introducing FedMR for multimodal data fusion on different FedRec methods under varying client sampling ratios, we conduct experiments on the five FedRec methods using the KU dataset. Fig. 8 presents the results for HR@50 and NDCG@50 as the client sampling ratio changes. Overall, as the client sampling ratio increases, most methods show improvements in HR@50 and NDCG@50, indicating that more client participation helps enhance the personalization of recommendations. However, the performance improvement varies across different methods. For PFedRec, the increase in HR@50 and NDCG@50 after introducing FedMR is relatively small, possibly due to its lower sensitivity to the number of participating clients. This suggests that PFedRec may rely more on the quality of individual client data rather than the number of clients. As a result, the advantages of FedMR’s multimodal fusion are not fully realized when the client count is low. Additionally, when the client sampling ratio is low, the model’s performance becomes unstable, indicating that insufficient data affects the stability of feature learning. This is especially true in multimodal scenarios, where a lack of sufficient modality information weakens the fusion effect, reducing recommendation accuracy.
Table 6. Performance comparison on the KU dataset. The table reports HR@50 and NDCG@50 for several FedRecs. Percentages indicate the performance drop relative to the full-feature baseline.
Method Metric FedMR w/o V w/ rnd V w/o C w/ rnd C w/o D w/ rnd D
FedAvg HR@50 0.5147 0.4069 (20.94% $\downarrow$) 0.3922 (23.80% $\downarrow$) 0.3725 (27.63% $\downarrow$) 0.3676 (28.58% $\downarrow$) 0.5094 (1.03% $\downarrow$) 0.4520 (12.18% $\downarrow$)
NDCG@50 0.1491 0.1391 (6.71% $\downarrow$) 0.1303 (12.61% $\downarrow$) 0.1287 (13.68% $\downarrow$) 0.1240 (16.83% $\downarrow$) 0.1407 (5.66% $\downarrow$) 0.1375 (7.78% $\downarrow$)
FCF HR@50 0.5539 0.2843 (48.67% $\downarrow$) 0.2892 (47.79% $\downarrow$) 0.4216 (23.89% $\downarrow$) 0.2451 (55.75% $\downarrow$) 0.2353 (57.52% $\downarrow$) 0.2206 (60.17% $\downarrow$)
NDCG@50 0.1597 0.0748 (53.16% $\downarrow$) 0.0783 (50.97% $\downarrow$) 0.0965 (39.57% $\downarrow$) 0.0629 (60.61% $\downarrow$) 0.0578 (63.81% $\downarrow$) 0.0538 (66.31% $\downarrow$)
FedNCF HR@50 0.1961 0.1569 (19.99% $\downarrow$) 0.1667 (14.99% $\downarrow$) 0.1863 (5.00% $\downarrow$) 0.1618 (17.49% $\downarrow$) 0.1602 (18.31% $\downarrow$) 0.1569 (19.99% $\downarrow$)
NDCG@50 0.0609 0.0431 (29.23% $\downarrow$) 0.0489 (19.70% $\downarrow$) 0.0541 (11.17% $\downarrow$) 0.0461 (24.30% $\downarrow$) 0.0538 (11.66% $\downarrow$) 0.0458 (24.79% $\downarrow$)
PFedRec HR@50 0.2157 0.1990 (7.74% $\downarrow$) 0.1647 (23.64% $\downarrow$) 0.1941 (10.01% $\downarrow$) 0.1696 (21.37% $\downarrow$) 0.1794 (16.83% $\downarrow$) 0.1647 (23.64% $\downarrow$)
NDCG@50 0.0616 0.0469 (23.86% $\downarrow$) 0.0468 (24.03% $\downarrow$) 0.0537 (12.82% $\downarrow$) 0.0445 (27.76% $\downarrow$) 0.0473 (23.21% $\downarrow$) 0.0484 (21.43% $\downarrow$)
FedRAP HR@50 0.3824 0.2647 (30.78% $\downarrow$) 0.2206 (42.31% $\downarrow$) 0.3235 (15.40% $\downarrow$) 0.2304 (39.75% $\downarrow$) 0.1863 (51.29% $\downarrow$) 0.1324 (65.38% $\downarrow$)
NDCG@50 0.1305 0.0762 (41.61% $\downarrow$) 0.0634 (51.42% $\downarrow$) 0.0714 (45.29% $\downarrow$) 0.0564 (56.78% $\downarrow$) 0.0490 (62.43% $\downarrow$) 0.0338 (74.10% $\downarrow$)
D.2.1. Ablation Study on Modality Contribution
We present the full modality contribution analysis, including results for noise replacement ($w/rnd$), in Table 6 owing to page limits.
We assess the individual contributions of visual ($\mathbf{V}$), textual ($\mathbf{C}$), and ID ($\mathbf{D}$) feature types within FedMR by evaluating performance upon either removing the feature type entirely ($w/o$) or replacing it with random noise ($w/rnd$). Table 6 presents these results on the KU dataset, revealing the distinct and complementary roles of each modality. The comparison between feature removal and noise replacement elucidates the nature of each modality’s contribution.
ID features consistently prove indispensable for personalization. Their removal ($w/o~{}D$) incurs substantial performance degradation across all methods. Notably, replacing ID features with random noise ($w/rnd~{}D$) often leads to even more severe performance drops, particularly for models with strong personalization components like FCF and FedRAP. This trend underscores that models heavily rely not just on the presence of an ID signal, but on the specific collaborative and structural information learned within the ID embeddings.
Visual features also demonstrate significant influence. Removing them ($w/o~{}V$) markedly hinders performance, especially for architectures known to leverage visual cues, such as FCF and FedRAP. Comparing this to noise replacement ($w/rnd~{}V$), we observe that substituting visual features with noise frequently results in equal or greater performance degradation than simple removal. This finding suggests that visual features provide meaningful, non-random signals that benefit recommendation, and replacing them with noise is often more detrimental than their absence, particularly for visually sensitive models.
Textual features provide valuable complementary semantic context. While their removal ($w/o~{}C$) impacts performance to a lesser extent than ID or visual features for some models, the effect is still notable across various methods. Replacing textual features with noise ($w/rnd~{}C$) generally causes a more significant performance decrease than simple removal, confirming that the semantic content captured by text offers a non-trivial contribution beyond just acting as a placeholder.
Collectively, these results highlight a clear interdependence among ID, visual, and textual information sources. The consistent performance drop upon removal or noise replacement confirms that each modality contributes uniquely. The often more detrimental effect of noise replacement compared to feature removal strongly indicates that the learned multimodal features contain meaningful signals crucial for performance. Furthermore, the degree of reliance on each modality, and the differential impact of removal versus noise replacement, varies markedly across the underlying FedRec architectures. This complex interplay and differential dependency motivate the need for adaptive fusion mechanisms. FedMR’s MFFM addresses this by dynamically balancing the contributions from these distinct and valuable information sources based on user context, enabling robust performance through effective information leverage.
Appendix E More Related Work
E.1. Federated Recommendation Systems
Federated recommendation systems (FedRec) aim to provide personalized recommendations while ensuring user privacy and data security (Yang et al., 2020). In recent years, FedRec has made significant progress in areas such as rating prediction (Lin et al., 2020a, b; Liang et al., 2021; Lin et al., 2022; Zhang et al., 2023d; Li et al., 2024b), top-k recommendation (Ammad-Ud-Din et al., 2019; Perifanis and Efraimidis, 2022; Luo et al., 2022; Zhang et al., 2023b; Dong et al., 2022; Li et al., 2022a; Qu et al., 2023), and addressing the cold-start problem (Liu et al., 2021b; Meihan et al., 2022; Wahab et al., 2022; Zhang et al., 2024c, d). However, most existing FedRec methods (Huang et al., 2020; Flanagan et al., 2021; Wu et al., 2022; Zhu et al., 2022; Imran et al., 2023; Yuan et al., 2023a, 2024) rely solely on ID embeddings to model item representations, overlooking the rich multimodal information (e.g., titles and covers) that items may have. AMMFRS (Feng et al., 2024) and FedMMR (Li et al., 2024a) apply multimodal learning (Baltrušaitis et al., 2018) to FedRec. AMMFRS uses an attention mechanism (Niu et al., 2021) to weight and fuse multimodal data, emphasizing the contribution of different features. In contrast, FedMMR assigns the computationally heavy multimodal learning tasks to the server and introduces a secure cross-client convolution strategy to capture high-order interaction information using secret sharing techniques while preserving privacy. Unlike methods that focus only on personalizing user information, PFedRec (Zhang et al., 2023c) proposes a dual personalization mechanism that optimizes personalization for both users and items, integrating it into existing FedRec frameworks. Building on this, FedRAP (Li et al., 2023a) introduces additive personalization, which refines the modeling of user and item features. FedDAE (Li et al., 2024c) employs a dual-encoder variational autoencoder to adaptively fuse shared and personalized item knowledge through a personalized gating network. While these methods achieve personalized item representations, they still mainly rely on ID embeddings, failing to fully utilize the rich multimodal information of items.
E.2. Multimodal Recommendation Systems
Unlike traditional Recommendation Systems (RecSys) that rely primarily on user-item interaction data (such as clicks or ratings) to capture user behavior (Yu et al., 2023), Multimodal Recommendation systems (MMRec) aim to improve the accuracy of recommendations and enhance user experience by utilizing multiple modalities of data , such as text, images, audio, and video (Zhou et al., 2023b). As unique identifiers for items, item IDs not only contain semantic information about the items but also capture the characteristics of neighboring items (Yuan et al., 2023b; Liu et al., 2023). Therefore, in MMRec, item IDs are often used as auxiliary information and combined with multimodal data to enhance item representations(Wang et al., 2024; Zhang et al., 2024e). This integration helps capture a more comprehensive understanding of users’ interests and preferences, ultimately improving the quality of recommendations. A typical MMRec training process generally involves feature extraction and modality fusion (He and McAuley, 2016; Wei et al., 2019; Wang et al., 2021; Tao et al., 2022). Feature extraction (Guo et al., 2019) aims to obtain representative features from multiple modalities to better capture the attributes and preferences of users and items. Modality fusion (Gao et al., 2020) focuses on effectively integrating data from different modalities to enhance the overall performance of the MMRec and the relevance of its results. The recommendation models (Tao et al., 2020; Wu et al., 2021; Zhou and Shen, 2023) combine multimodal data with user-item interaction data to predict user preferences and provide personalized recommendations.
E.3. Foundation Models
Foundation Models (FM) are often pre-trained on large-scale unlabeled datasets for general purposes, that can be adapted to a variety of downstream tasks through appropriate fine-tuning, such as task-specific training (Bommasani et al., 2021). This makes them highly generalizable and adaptable across different applications. They are then adapted to specific recommendation tasks through methods such as representation transfer (Du et al., 2020; Liu et al., 2021a, 2022; Wei et al., 2023), model fine-tuning (Zhang et al., 2021; Li et al., 2022b, 2023b; Sun et al., 2023; Wang et al., 2023), adapter tuning (Hou et al., 2022; Bian et al., 2023; Geng et al., 2023; Fu et al., 2024), and prompt tuning (Shin et al., 2020; Li and Liang, 2021; Zhou et al., 2022; Zhang and Wang, 2023; Xu et al., 2024). Some studies have explored the use of FM to enhance the content-awareness of MMRec, improving content understanding and generation by leveraging alignment and generation techniques across multiple modalities. CLIP (Radford et al., 2021a; Wu et al., 2024) uses contrastive learning (Le-Khac et al., 2020) between language and images to obtain multimodal representations, capturing the matching relationships between users and items. M6-Rec (Cui et al., 2022) builds a recommendation system based on large-scale multimodal pre-trained models specifically designed for the e-commerce domain. P5 (Geng et al., 2022) transforms the recommendation task into a natural language processing task, generating personalized recommendations through prompts. VIP5 (Geng et al., 2023) leverages pre-trained multimodal models to capture complementary information from visual and textual data, providing more personalized and relevant recommendations to users. FedPA (Zhang et al., 2024c) and TransFR (Zhang et al., 2024b) integrate FM with FedRec. FedPA introduces a personalized adapter to fine-tune the FM, allowing each client to learn individual user preferences in a lightweight manner. In contrast, TransFR leverages FMs to generate universal textual representations, replacing traditional ID-based item representations and enhancing the system’s adaptability to new domains and its ability to handle cold-start issues.
Appendix F Limitations and Future Work
While FedMR demonstrates promising results for privacy-preserving multimodal recommendation, key limitations suggest avenues for future research. The specific subfield of FedMMRec is in its early stages: while FedMR advances this area alongside recent works like AMMFRS (Feng et al., 2024) and FedMMR (Li et al., 2024a), establishing direct experimental comparisons to precisely gauge relative performance remains a key challenge for future work as the field evolves and common benchmarks emerge. Additionally, FedMR primarily utilizes visual and textual data via CLIP, future research should explore incorporating a broader range of relevant modalities, such as audio or temporal information from videos. Further investigations could also focus on enhancing the client-side MFFM personalization mechanism or integrating FedMR with advanced FL techniques to better address communication constraints or statistical heterogeneity.